### v1

# EvalAlign: Supervised Fine-Tuning Multimodal LLMs with Human-Aligned Data for Evaluating Text-to-Image Models

Abstract
The recent advancements in text-to-image generative models have been remarkable. Yet, the field suffers from a lack of evaluation metrics that accurately reflect the performance of these models, particularly lacking fine-grained metrics that can guide the optimization of the models. In this paper, we propose EvalAlign, a metric characterized by its accuracy, stability, and fine granularity. Our approach leverages the capabilities of Multimodal Large Language Models (MLLMs) pre-trained on extensive datasets. We develop evaluation protocols that focus on two key dimensions: image faithfulness and text-image alignment. Each protocol comprises a set of detailed, fine-grained instructions linked to specific scoring options, enabling precise manual scoring of the generated images. We Supervised Fine-Tune (SFT) the MLLM to align closely with human evaluative judgments, resulting in a robust evaluation model. Our comprehensive tests across 24 text-to-image generation models demonstrate that EvalAlign not only provides superior metric stability but also aligns more closely with human preferences than existing metrics, confirming its effectiveness and utility in model assessment.
1 Introduction
Text-to-image models, such as DALL·E series [38, 37, 3], Imagen [43], and Stable Diffusion [33], have significantly impacted various domains such as entertainment, design, and education, by enabling high-quality image generation. These technologies not only advance the field of text-to-image generation but also enhance related applications such as video generation [4, 60], image editing [46, 18, 59], and personalized image generation [13, 42]. Despite this prevalence, in the practical use of generative models, we still lack evaluation methods to accurately evaluate existing text-to-image models.
Existing benchmarks for text-to-image generation models are neither comprehensive nor accurate enough to meet current evaluation needs. The reasons for this include: (1) Limited model parameters: Current evaluation models have too few parameters, which restricts their ability to accurately represent images, leading to significant discrepancies compared to human evaluations. (2) Training data limitations: Some evaluation methods, such as Inception Score (IS) [44], Frechet Inception Distance (FID) [15], and CLIP Score [14], use models that have not been trained with synthesized images, which may introduce training bias and flaws the evaluation. (3) High annotation costs: Some methods, such as ImageReward [56], HPS [55] and HPS v2 [54], rely heavily on extensive human annotations, which significantly increases the cost of labeling. (4) Lack of detailed evaluation metric: The evaluation metrics do not provide fine-grained interpretability, preventing them from guiding model optimization effectively. (5) Computational inefficiency: The evaluation models require substantial computational resources, making them inefficient.
To address these issues, we propose EvalAlign, which offers low-cost, accurate, and efficient model evaluations while providing fine-grained, interpretable metrics. Specifically, we Supervised Fine-Tune (SFT) a Multimodal Large Language Model (MLLM) to align it with human annotations. Our focus is on two key evaluation aspects: image faithfulness and text-image alignment.
Due to its pre-training on large-scale datasets and a large number of model parameters, the MLLM demonstrates excellent understanding and generalization capabilities for images and instructions. This makes it possible to design evaluation instruction based on MLLM to evaluate text-to-image models. However, since the pre-trained datasets do not include model-generated images (such as distorted body structures and human hands) or evaluation-related text instructions, using MLLM directly for model evaluation does not yield optimal results. Therefore, we employ SFT on a small amount of manually annotated data to align MLLM with human annotations for text-to-image generation. In summary, our main contributions can be summarized as follows:
We introduce a dataset specifically designed to address the evaluation challenges of text-to-image models. This dataset, derived from multiple data sources, has been thoroughly cleaned and systematically annotated by human. It enables precise evaluation of text-image alignment and image faithfulness.
We propose the EvalAlign evaluation metric, which accurately aligns evaluation metrics with human preferences. This method is cost-effective in terms of annotation and training, computationally efficient, and provides interpretable metrics.
We conduct evaluations over 24 text-to-image models and compare EvalAlign with existing evaluation methods. Extensive experiments demonstrate that EvalAlign outperforms other methods in evaluating model performance.
2 Related Work
Evaluations and benchmarks of text-to-image generation. Despite the incredible progress achieved by text-to-image generation, evaluations and benchmarks in this area are far from flawless and contain critical limitations. For example, the most commonly used metrics, IS [44], FID [15], and CLIPScore [14] are broadly recognized as inaccurate for their inconsistency with human perception. To address, HPS series [55, 54], PickScore [20], and ImageReward [56] introduced human preference prior on image assessing to the benchmark, thereby allowing better correlation with image quality. However, with varying source and size of training data, these methods merely score the evaluated images in a coarse and general way, which cannot serve as an indication for model evolution. Meanwhile, HEIM [21] combined automatic and human evaluation and holistically evaluated text-to-image generation in 12 aspects, such as alignment, toxicity, and so on. As a consequence, HEIM relies heavily on human labour, limiting its application within budget-limited research groups severely. [30] standardized the protocol and settings of human evaluation, ensuring its verifiable and reproducible. There are also some works bear a resemble with us. For instance, TIFA [16], Gecko [53] and LLMScore [27] also formulate the evaluation as a set of visual question answering procedure and use LLMs as evaluation models. However, while they all mainly focus on text-image alignment, our approach takes both text-image alignment and image faithfulness into consideration. Moreover, the evaluation of LLMScore requires an object detection stage, which introduces significantly extra inference latency to the evaluation pipeline.
As illustrated in Table 1, existing text-to-image evaluation methods contains various limitations, making them incapable to serve as a fine-grained, comprehensive, and human-preference aligned automatic benchmark. While our work fills in this gap economically, and can be employed to indicate evolution direction and support thorough analysis of text-to-image generation models.
| Method | Venue | Benchmark Feature | Dataset Size | Evaluation Aspect | |||||
| Human-aligned | Fine-grained | Automatic | P | I | A | Faithfulness | Alignment | ||
| Inception Score [44] | NeurIPS 2016 | ✗ | ✗ | ✓ | – | 1.3M | – | ✓ | ✗ |
| FID [15] | NeurIPS 2017 | ✗ | ✗ | ✓ | – | 1.3M | – | ✓ | ✗ |
| CLIP-score [14] | EMNLP 2021 | ✗ | ✗ | ✓ | 400M | 400M | – | ✗ | ✓ |
| HPS [55] | ICCV 2023 | ✓ | ✗ | ✓ | 25K | 98K | 25K | – | – |
| TIFA [16] | ICCV 2023 | ✓ | ✓ | ✓ | 4K | – | 25K | ✗ | ✓ |
| TVRHE [30] | CVPR 2023 | ✓ | ✗ | ✗ | – | – | – | ✓ | ✗ |
| ImageReward [56] | NeurIPS 2023 | ✓ | ✗ | ✓ | 8.8K | 68K | 137K | – | – |
| PickScore [20] | NeurIPS 2023 | ✓ | ✗ | ✓ | 35K | 1M | 500K | – | – |
| HPS v2 [54] | arXiv 2023 | ✓ | ✗ | ✓ | 107K | 430K | 645K | – | – |
| HEIM [21] | NeurIPS 2023 | ✓ | ✓ | ✗ | – | – | – | ✓ | ✓ |
| Gecko [53] | arXiv 2024 | ✓ | ✓ | ✓ | 2K | – | 108K | ✗ | ✓ |
| LLMScore [27] | arXiv 2024 | ✓ | ✓ | ✓ | – | – | – | ✗ | ✓ |
| EvalAlign (ours) | – | ✓ | ✓ | ✓ | 3K | 21K | 132K | ✓ | ✓ |
Multimodal Large Language Models (MLLMs). Pretrained on massive text-only and image-text data, MLLMs have exhibited exceptional image-text joint understanding and generalization abilities, facilitating a large spectrum of downstream applications. Among the works major in MLLMs, LLaVA [26, 24] and MiniGPT4 [61, 5] proposed to conduct visual instruction tuning during SFT, so that MLLMs can be easily aligned to human preference and precisely answer fine-grained questions on visual content. Meanwhile, Video-LLaMA [58] and VideoChat [22] utilized MLLMs for video understanding. VILA [23] quantitatively proved that involving text-only instruction-tuning data during SFT can further ameliorate model performance on text-only and multimodal downstream tasks. LLaVA-NeXT [25] extracted visual tokens for both the resized input image and the segmented sub-images to provide more detailed visual information for MLLMs, achieving significant performance bonus on tasks with high-resolution input images.
In this paper, we manually annotate a minimal amount of high-quality visual instruction-tuning data to adapt MLLMs to perform human preference aligned and fine-grained evaluation. Owing to the generalization and multimodal understanding abilities of MLLMs, our experiments demonstrated that the finetuned MLLMs can accurately assess text-to-image generation models by following instructions regarding various fine-grained criteria of synthesised image.
3 EvalAlign Dataset Construction
To train, validate and test the effectiveness of our evaluation models, we build EvalAlign benchmark. Specifically, EvalAlign dataset is a meticulously annotated collection featuring fine-grained annotations for images generated from text prompts. This dataset comprises 21k images, each accompanied by detailed instructions. The compilation process for the EvalAlign Dataset encompasses prompt collection, image generation, and precise instruction-based annotation.
3.1 Prompts and Images Collection
Prompt collection. To assess the capabilities of our model in terms of text-image alignment and image faithfulness, we collect, filter, and clean prompts from existing evaluation datasets and generated prompts based on LLM. These prompts encompass a diverse range from real-world user prompts, prompts generated through rule-based templates with LLM, to manually crafted prompts. Specifically, the utilized prompts are soureced from HPS [55], HRS-Bench [2], HPSv2 [54], TIFA [16], DSG [7], T2I-Comp [17], Winoground [50], DALL-EVAL [8], DiffusionDB [52], PartiPrompts [57], DrawBench [43], and JourneryDB [49].
Prompt curation. The prompts are collected for facilitating fine-grained and comprehensive evaluation on text-to-image models in terms of image faithfulness and text-image alignment. Considering some of the collected prompt are unsuitable for these two evaluation tasks, we filter the collected prompts to ensure their quantity, quality and diversity. For the image faithfulness evaluation task, we prioritize prompts related to human, animals, and other tangible objects, as prompts depicting sci-fi scenarios are less suitable for this type of assessment. Our filtering process initially selects prompts that describe human, animals, and other real objects. After deduplicating these prompts, we carefully selected 1,500 distinct prompts with varying topic, background and style for further evaluation. The selected prompts encompass 10k subjects across 15 categories. For the text-image alignment task, we refined our selection based on descriptions of style, color, quantity, and spatial relationships in the prompts. Only those prompts containing relevant descriptions and exceeding 15 words in length were considered, culminating in a final set of 1,500 prompts for evaluation.
Image generation. To train and evaluate the MLLM, we use a diverse set of images generated by various models using the aforementioned prompts, allowing for detailed human annotation. For each prompt, multiple images are generated across different models. The dataset is enriched by images synthesized by models with varying architectures and scales, enhancing the diversity essential for a comprehensive evaluation.
This variety not only tests the MLLM generalization capabilities but also aids in developing a model with broader applicability. The training dataset incorporates images from 8 models, whereas the test dataset spans 24 models. The inclusion of 16 unseen models in the test set is crucial for evaluating the MLLM to generalize beyond its training data. For detailed information on the inference setting of each model, please refer to the supplementary material. This structured approach ensures a robust framework for training and validating the MLLM, positioning it as a versatile and adaptive tool in the field of image generation from textual descriptions.
3.2 Data Annotation
Prompt annotation. For text prompts focused on text-image alignment, we begin by annotating the entities and their attributes within the text, as illustrated in Figure 1. Our annotators extract the entities mentioned in the prompts and label each entity with corresponding attributes, including quantity, color, spatial relationships, and actions. During the annotation, we also ask the annotators annotate the overall style of the image if described in the corresponding prompt and report prompts that contain toxic and NSFW content. These high-quality and detailed annotations facilitate the subsequent SFT training and evaluation of the MLLM. This meticulous annotation procedure ensures that the MLLM can accurately align and respond to the nuanced details specified in the prompts, enhancing both the training process and the model’s performance in generating images that faithfully reflect the described attributes and style.
Image annotation. The images generated by text-to-image models often present challenges such as occluded human body parts, which can impede the effectiveness of SFT training and evaluation of the MLLM. Therefore, to address these challenges and enhance the model’s training and evaluative capabilities, specific annotations are applied to all images depicting human and animals. These annotations include: presence of human or animal faces; visibility of hands; visibility of limbs. By implementing these annotations, we ensure that the MLLM can more effectively learn from and assess the completeness and faithfulness of the generated images. This structured approach to annotation not only aids in identifying common generation errors but also optimizes the model’s ability to generate more accurate and realistic images, thereby improving both training outcomes and the model’s overall performance in generating coherent and contextually appropriate visual content.
Instruction-Fine-tuning data annotation. To align the MLLM with human preference prior on fine-grained image assessing, we can train the model on a minimal amount of fine-grained human preference data through SFT training. As a consequence, we devise two sets of questions, each is concentrated on a specific fine-grained aspect of image faithfulness and image-text alignment, and ask human annotators to answer these questions to acquire the fine-grained human preference data. To aid the human annotators to understand the meaning of each question and its available answer options, thereby ensuring high annotation quality, we employ a thorough and comprehensive procedure of annotation preparation. First, we write a detailed annotation guideline and conduct a training for the annotators to explain the annotation guideline and answer their questions about the annotation (see supplementary materials for the annotation guideline). Then, we conduct a multi-turn trial annotation on another 50 synthesized images. After each trial, we calculate the Cohen’s kappa coefficient and interpret annotation guidelines for our annotators. In total, we conduct nine turns of trial annotation, and in the last turn of the trial, the Cohen’s kappa coefficient of our annotators reaches , indicating high inter-annotator reliability and high annotation quality. More results about the trial annotation will be included in the supplementary materials.
After completing the aforementioned preparations, we delegate the images filtered during image annotation to 10 annotators and ask them to complete the annotation just as how they did in the trial annotation. Furthermore, during the whole annotation procedure, four text-to-image generation experts conduct random sampling quality inspection on the present annotated results, causing a second and a third re-annotation on 423 and 112 inspection-failed samples. Overall, owing to the valuable work of our human annotators and our fastidious annotation procedure, we get quality-sufficient instruction-tuning data required for the SFT training of the MLLM. More details of the annotation procedure will be introduced in the supplementary files.
3.3 Dataset Statistics
To summarize, we generate 24k images from 3k prompts based on 8 text-to-image models, which includes DeepFloyd IF [1], SD15 [40], LCM [28], SD21 [40], SDXL [33], Wuerstchen [32], Pixart [6], and SDXL-Turbo [48]. After data filtering, 4.5k images were selected as annotation data for task of text-image alignment. Subsequently, these images were carefully annotated to generate 13.5k text-image pairs, where 11.4k were used to the training dataset and 2.1k to the validation dataset. For the image faithfulness task, we select 12k images for annotate, producing 36k text-image pairs, with 30k are used to the training dataset and 6.2k to the validation dataset. Additionally, we employed 24 text-to-image models to generate 2.4k images from 100 prompts. After annotation, these images were used as testing dataset. Figure 3 and Figure 3 show the distribution of objects in different categories within our prompts, demonstrating the diversity and balance of our prompts.
4 Training and Evaluation Methods
4.1 Supervised Fine-Tuning the MLLM
As we mentioned above, we use MLLMs as the evaluation models and let it to answer a set of carefully-designed instructions, thereby achieving quantitative measurement of fine-grained text-to-image generation skills. Due to training bias, zero-shot MLLMs perform poorly when it comes to fine-grained evaluation on synthesized images, particularly in term of image faithfulness. To address, we conduct SFT training on the annotated data to align the MLLM with human preference prior on fine-grained image assessing. Formally, the SFT training sample can be denoted as a triplet, i.e., , where denotes the question (or the instruction), the multimodal input and the answer, respectively. Specifically, is a question about a fine-grained aspect of the generated image as well as the available answer options, is mainly the image and necessary textual description, while is the specific answer chosen from the given options and annotated by human. Notably, we expand each given option to a more detailed and descriptive version, benefiting the SFT performance by allowing the MLLM to understand the meaning of each option. During SFT training, the optimization objective is the original autoregressive loss function utilized to train LLMs, but calculated only on . Specifically, the loss function can be formulated as follows, where is the length of the ground truth answer:
4.2 Evaluation and Metrics
During inference, the MLLM, parameterized by , is required to generate the appropriate response , given the specific question and the multimodal input in an autoregressive way:
This autoregressive generation procedure is considered complete once the model generates a EOS token or the generated response exceeds the preset maximum generation length. After the generation, we use rule-based filtering and regular expression to extract the option chosen by the MLLM, where each option has an exclusive predefined score to quantitatively measure a fine-grained skill specified by the question :
where represents the procedure of option extraction and score searching. We devise two holistic and detailed question sets, and , that encompass every aspect of image faithfulness and text-image alignment, respectively. Consequently, Our metric EvalAlign can be defined by simply summing the scores of the two question sets:
where and indicates the image faithfulness score and image-text alignment score evaluated by EvalAlign.
4.3 Implementation Details
We leverage the LLaVA-NeXT model and applied the LoRA fine-tuning method to align it with the training dataset as described previously. The entire training process was conducted on NVIDIA A100 GPUs with hours, using a learning rate of 510-5. Additionally, we restricted the LoRA fine-tuning to the Q and K weights of the attention module, as extending the fine-tuning to the ViT and projection modules was found to cause overfitting. Importantly, our method does not necessitate any modifications to the MLLM, rendering it compatible with any MLLM model. For the ablation study, we evaluated the LLaVA-NeXT 13B model on the validation dataset. In the final experiment, we apply supervised fine-tuning to the LLaVA-NeXT 34B model on the testing dataset and compare against different evaluation metrics.
5 Experimental Results
In this section, we show results of EvalAlign on benchmarking 24 text-to-image generative models. We also compare with popular text-to-image metrics and frameworks, including Inception Score (IS) [44], Frechet Inception Distance (FID) [15], CLIP-score [14], and HPS v2 [54]. Please refer to supplementary materials for detailed results and analyses.
5.1 Main Results
| Model | Human | EvalAlign |
| PixArt XL2 1024 MS [6] | 2.2848 | 1.6415 |
| Dreamlike Photoreal v2.0 [11] | 2.0070 | 1.4522 |
| SDXL Refiner v1.0 [47] | 1.9229 | 1.6072 |
| SDXL v1.0 [33] | 1.8136 | 1.4675 |
| Wuerstchen [32] | 1.7837 | 1.4279 |
| LCM SDXL [28] | 1.6910 | 1.3391 |
| Openjourney [34] | 1.6667 | 1.175010 |
| Safe SD MAX [31] | 1.6491 | 1.2175 |
| LCM LORA SDXL [28] | 1.6387 | 1.3833 |
| Safe SD STRONG [31] | 1.630810 | 1.146611 |
| Safe SD MEDIUM [31] | 1.627511 | 1.129815 |
| Safe SD WEAK [31] | 1.607812 | 1.118817 |
| SD v2.1 [40] | 1.552413 | 1.109418 |
| SD v2.0 [40] | 1.527714 | 1.130014 |
| Openjourney v2 [35] | 1.500015 | 0.995620 |
| Redshift diffusion [39] | 1.473316 | 1.138212 |
| Dreamlike Diffusion v1.0 [10] | 1.465217 | 1.2052 |
| SD v1.5 [40] | 1.441718 | 1.136213 |
| IF-I-XL v1.0 [1] | 1.380819 | 0.922122 |
| SD v1.4 [40] | 1.359220 | 0.951121 |
| Vintedois Diffusion v0.1 [51] | 1.356221 | 1.079719 |
| IF-I-L v1.0 [1] | 1.263522 | 0.881423 |
| MultiFusion [29] | 1.237223 | 1.129816 |
| IF-I-M v1.0 [1] | 1.013524 | 0.792824 |
| Model | Human | EvalAlign |
| IF-I-XL v1.0 [1] | 5.4500 | 5.5300 |
| IF-I-L v1.0 [1] | 5.2300 | 5.4500 |
| SDXL Refiner v1.0 [47] | 5.2100 | 5.4000 |
| LCM SDXL [28] | 5.1800 | 5.3300 |
| PixArt XL2 1024 MS [6] | 5.1100 | 5.3100 |
| IF-I-M v1.0 [1] | 5.0800 | 5.2200 |
| LCM LORA SDXL [28] | 5.0600 | 5.2700 |
| SDXL v1.0 [33] | 5.0300 | 5.3500 |
| Wuerstchen [32] | 4.8700 | 5.1700 |
| Openjourney [34] | 4.830010 | 4.920015 |
| SD v2.1 [40] | 4.800011 | 5.070011 |
| MultiFusion [29] | 4.680012 | 4.800018 |
| Dreamlike Diffusion v1.0 [10] | 4.660013 | 5.150010 |
| SD v2.0 [40] | 4.640014 | 5.010012 |
| Vintedois Diffusion v0.1 [51] | 4.620015 | 4.950014 |
| Safe SD STRONG [31] | 4.600016 | 4.830017 |
| Dreamlike Photoreal v2.0 [11] | 4.560017 | 4.980013 |
| Safe SD WEAK [31] | 4.530018 | 4.710020 |
| SD v1.4 [40] | 4.520019 | 4.760019 |
| SD v1.5 [40] | 4.450020 | 4.900016 |
| Safe SD MEDIUM [31] | 4.400021 | 4.560024 |
| Redshift diffusion [39] | 4.350022 | 4.670021 |
| Safe SD MAX [31] | 4.310023 | 4.590023 |
| Openjourney v2 [35] | 4.150024 | 4.650022 |
Evaluation on image faithfullness. We conducted the image faithfullness evaluation on the testing dataset, as detailed in Table 3. Our evaluation metric, EvalAlign, demonstrates close alignment with human preferences Specifically, the rankings of the top and bottom 10 models by both EvalAlign and human evaluation scores show remarkable consistency, confirming that our metric closely mirrors human evaluation. Importantly, most of the images from the 24 generative models were not present during training, showcasing the robust generalization capability of our evaluation method.
Evaluation on text-image alignment. The evaluation of text-image alignment on the testing dataset was conducted follow the same manner of image faithfullness evaluation. Table 3 reveals that the rankings based on the EvalAlign metric and human evaluation metric for the 24 models are generally consistent. This consistency likely stems from the proximity of MLLM’s pre-training tasks to the text-image alignment evaluation tasks, resulting in better performance on text-image alignment evaluation task.
5.2 Ablations and Analyses of EvalAlign
Results on different prompt categories. SFT significantly enhances the performance of MLLMs on evaluation tasks. We conducted experiments comparing the LLava-Next 13B model with and without SFT. As shown in Table 5 and Table 5, the results demonstrate that SFT considerably improves performance across all prompt categories in both image faithfulness and text-to-image alignment, closely aligning the MLLM’s predictions with human evaluations. Note that Table 5 shows that the baseline method without SFT performs poorly on the image faithfulness task, which suggests the necessity of SFT.
| Method | Body | Hand | Face | Object | Common |
| Human | 1.6701 | 1.0278 | 1.4107 | 2.2968 | 1.0637 |
| Baseline | 3.9950 | 3.9932 | 3.9867 | 2.6734 | 3.3476 |
| EvalAlign | 1.7305 | 0.9490 | 1.4393 | 2.3565 | 1.0903 |
| Method | Object | Count | Color | Style | Spatial | Action |
| Human | 1.6947 | 1.2032 | 1.8551 | 1.9796 | 1.5608 | 1.8015 |
| Baseline | 1.5602 | 1.0742 | 1.9275 | 1.1837 | 1.4118 | 1.1838 |
| EvalAlign | 1.6807 | 1.2516 | 1.8696 | 1.9592 | 1.5882 | 1.8382 |
Effect of dataset size for vision-language model training. In order to explore the effects of data size, we train model on image faithfulness evaluation task under three different data size: 200, 500 and 800. As illustrated in Table 6, performance generally enhances with more training data. Notably, training with just 500 data nearly maximizes accuracy, with further increases to 800 data yielding only slight improvements. This result suggests that our method requires only a small amount of annotated data to achieve good performance, highlighting the cost-effectiveness of our approach.
| Method | Data Size | SDXL | Pixart | Wuerstchen | SDXL-Turbo | IF | SD v1.5 | SD v2.1 | LCM |
| Human | – | 2.1044 | 1.8606 | 1.7839 | 1.3854 | 1.3822 | 1.3818 | 1.1766 | 1.0066 |
| EvalAlign | 200 | 1.7443 | 1.8898 | 1.9278 | 1.1261 | 1.2977 | 1.5254 | 1.4309 | 1.1204 |
| 500 | 1.8890 | 1.9161 | 1.8586 | 1.2141 | 1.3109 | 1.3926 | 1.3815 | 0.9485 | |
| 800 | 2.0443 | 1.9199 | 1.8012 | 1.3353 | 1.296 | 1.4702 | 1.3221 | 1.0305 |
Effect of model size. Transformers have demonstrated a strong capacity to scale effectively across a variety of tasks, as noted in key studies [9, 36]. In Table 7, we illustrate the advantages of scaling up the MLLM for image faithfulness evaluation. Specifically, increasing the MLLM model size from 7B to 34B results in notable improvements in performance.
| Method | Model Size | SDXL | Pixart | Wuerstchen | SDXL-Turbo | IF | SD v1.5 | SD v2.1 | LCM |
| Human | – | 2.1044 | 1.8606 | 1.7839 | 1.3854 | 1.3822 | 1.3818 | 1.1766 | 1.0066 |
| EvalAlign | 7B | 1.9959 | 1.8615 | 1.8228 | 1.1708 | 1.2704 | 1.4031 | 1.3063 | 1.0145 |
| 13B | 2.0443 | 1.9199 | 1.8012 | 1.3353 | 1.2960 | 1.4702 | 1.3221 | 1.0305 | |
| 34B | 2.1131 | 1.8621 | 1.8083 | 1.3906 | 1.3076 | 1.3921 | 1.2037 | 1.0143 |
5.3 Comparison with Existing Evaluation Methods
SFT with human-aligned data outperforms vanilla MLLMs. To validate the effectiveness of the MLLM after SFT, we use vanilla LLaVA-NeXT 13B as the baseline model for comparison. As shown in Table 5 and Table 5, the results of vanilla model suggest some correlations with human-annotated data. However, the alignment of the vanilla MLLM is relatively low due to the absence of images generated by model (such as distorted bodies and hands images) and issues related to evaluation in the MLLM’s pre-training dataset. After applying SFT on the LLaVA-Next 13B model using human annotated data, the model’s predictions on various fine-grained evaluation metrics are almost align to the human-annotated data and significantly surpass the evaluation results of all MLLM models that were not fine-tuned. This experimental results confirms that our SFT training enables the MLLM to be successfully applied to the task of evaluating text-to-image models.
| Method | Faithfulness | Alignment | ||
| Kendall | Pearson | Kendall | Pearson | |
| CLIP-score | 0.1304 | 0.1765 | 0.6956 | 0.8800 |
| HPSv2 | 0.4203 | 0.5626 | 0.5217 | 0.7113 |
| EvalAlign | 0.7464 | 0.8730 | 0.8043 | 0.9356 |
Comparison with state-of-the-art methods. To verify the human preference alignment of our model, especially when compared with other baseline methods, we calculate Kendall rank [19] and Pearson [12] correlation coefficient and summarize the results in Table 8. This comparison experiment encompasses images generated by 24 text-to-image models, covering an extensive range of synthesized image distribution. As can be concluded, compared with baseline methods, EvalAlign achieves significant higher alignment with fine-grained human preference on image faithfulness and image-text consistency, showcasing robust generalization ability. Although HPS v2 roughly aligns with human preference in some extent, the relative small model capacity and coarse ranking training limits its generalization to the fine-grained annotated data. Besides, since CLIP-s only cares the CLIP similarity of the generated image and its corresponding prompt, it behaves poorly in image faithfulness evaluation. The Per-question alignment and the final benchmark leaderboard of EvalAlign will be introduced in the supplementary materials.
6 Conclusion and Discussion
In this work, we design an economic and efficient evaluation method that offers high accuracy, strong generalization capabilities, and provides fine-grained, interpretable metrics. We develop a comprehensive data annotation and cleaning process tailored for evaluation tasks, and established the EvalAlign benchmark for training and evaluating models on supervised fine-tuning tasks for MLLMs. Experimental results across 24 text-to-image models demonstrate that our evaluation metrics surpass the accuracy of all the state-of-art evaluation method. Additionally, we conduct a detailed empirical study on how MLLMs can be applied to model evaluation tasks. There are still many opportunities for further advancements and expansions based on our EvalAlign. We hope that our work can inspire and facilitate future research in this field.
Checklist
The checklist follows the references. Please read the checklist guidelines carefully for information on how to answer these questions. For each question, change the default [TODO] to [Yes] , [No] , or [N/A] . You are strongly encouraged to include a justification to your answer, either by referencing the appropriate section of your paper or providing a brief inline description. For example:
Did you include the license to the code and datasets? [Yes] See Section X.
Did you include the license to the code and datasets? [No] The code and the data are proprietary.
Did you include the license to the code and datasets? [N/A]
Please do not modify the questions and only use the provided macros for your answers. Note that the Checklist section does not count towards the page limit. In your paper, please delete this instructions block and only keep the Checklist section heading above along with the questions/answers below.
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes] The discussion of limitation is included in the supplementary files.
Did you discuss any potential negative societal impacts of your work? [Yes] The discussion of negative societal impacts is recognised as part of the limitation, and will be included in the supplementary files.
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] The instruction is given in the project page link of the paper.
- (b)
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] All the experiments are reproduced by our members with different random seed.
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See 4.3.
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes]
Did you mention the license of the assets? [Yes] They will be mentioned in the Github repository if available.
Did you include any new assets either in the supplemental material or as a URL? [Yes]
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes]
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See 3.2.
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] They will be included in supplementary files.
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] The hourly wage for the works is 30 CNY/hour for every annotator.
Appendix A Limitations
Multimodal LLMs. Since EvalAlign evaluation models are fine-tuned MLLMs, they also suffer from multimodal hallucination, where models may generate content that seems plausible but actually incorrect or fabricated, and cannot be inferred from the input images and texts. Moreover, due to the possible harmful content in the pretraining data of the utilized base MLLMs, the model may inherit these biases and generate inappropriate response. Although we carefully curate the SFT training data of the EvalAlign evaluation models, the problems of hallucination and biased pre-training is alleviated but not fully addressed. Other than the these issues, EvalAlign evaluation models also suffer from opacity and interpretability, context limitation, as well as sensitivity to input formatting, like most multimodal LLMs.
Human Annotations. Human annotation is naturally subjective and influenced by individual perspectives, biases, and preferences. During the annotation, annotators can make mistakes, leading to incorrect or noisy labels. Regarding these challenges, we conduct 9 rounds of trial annotation and 2 rounds of random sampling quality inspection to ensure the inter-annotator consistency and overall annotation quality. We also design easy-to-understand annotation guidelines, instructions and platform to lower the annotation difficulty and benefit the annotation accuracy. Despite all these efforts, conducting human annotation with different annotators, user interface and annotation guidelines may lead to different result, making our annotation somewhat limited. Furthermore, human annotation can be time-consuming and resource-intensive, limiting the scale at which we can afford.
Appendix B Ethical and Social Impacts
Similar with other MLLMs, EvalAlign may potentially generate responses contain offensive, inappropriate, or harmful content. Since the base MLLMs of EvalAlign are pretrained on large datasets scraped from the web that might contain private information and harmful content, they may inadvertently generate or expose sensitive information, raising ethical and privacy concerns. MLLMs are also susceptible to adversarial attacks, where inputs are intentionally crafted to deceive the model. This vulnerability can be exploited to manipulate model outputs, posing security and ethic risks. To alleviate these safety limitation, we create dedicated evaluation sets for bias detection and mitigation, and conducted adversarial testing through hours of redteaming. Besides, EvalAlign is designed for fine-grained, human-aligned automatic text-to-image evaluations. We believe that with appropriate use, it could provide users with interesting experiences for detailed synthesized image evaluation, and inspires more appealing research works about text-to-image generation.
Appendix C Annotation Details
Before performing the final human annotation, we made a series of efforts to guarantee its quantity, quality and efficiency. To begin with, we select appropriate candidates to perform the annotation and hold a training meeting for them. Then, we design a user-friendly user interface and a comprehensive annotation procedure. We write a detailed annotation guidelines to explain every aspect and precaution of the annotation. As mentioned above, we conduct 9 rounds of trial annotation on another 50 synthesized images and 2 turns of random sampling quality inspection to further ensure inter-annotator consistency and annotation accuracy.
Annotator selection. The accuracy and reliability of the annotated data depend heavily on the capabilities of the human annotators involved in the annotation process. As a consequence, at the beginning of the annotation, We first conduct annotator selection to build an appropriate and unbiased annotation team, and train this annotation team with our meticulously prepared annotation guidelines. For annotator selection, we let the candidates to accomplish a test concentrating on 10 factors, domain expertise, resistance to visually disturbing content, attention to detail, communication skills, reliability, cultural and linguistic competence, technical skills, ethical considerations, aesthetic cognition, and motivation. Notably, since the evaluated models may generate images with uncomfortable and inappropriate visual content, the candidates are notified with this inconvenience before the test. Only those agreed with this inconvenience are eligible to participate in the test, and they are welcome to withdraw at any time if they choose to do so. Based on the test results and candidate backgrounds, We try our best to ensure that the selected annotators are well-balanced in background and have a generally competitive abilities of the 10 mentioned factors. To summarize, our annotation team includes 10 annotators carefully selected from 29 candidates, 5 males and 5 females, all have a bachelor’s degree. We interview the annotators and ensure they are adequate for the annotation.
Annotation training and guidelines. After the selection, we conduct a training meeting over our comprehensive user guidelines to make the annotation team aware of our purpose and standard. During the training meeting, we explain the purpose, precaution, standard, workload and wage of the annotation. Besides, we have formally informed the annotators that the human annotation is solely for research purposes, and the data they have annotated may potentially be released to the public in the future. We, and the annotators reached consensus on the standard, workload, wage and intended usage of the annotated data. The rules for recognising image faithfulness and text-image are universal, and thus each individual’s standards should not differ significantly. As a consequence, we annotate a few samples using our meticulously developed annotation platform for the annotators to ensure inter-annotator consistency. The overall snapshot of the developed annotation paltform is exhibited in fig. 4. With this training, we also equip the annotators with necessary knowledge for unbiased detailed human evaluation on image faithfulness and text-image alignment. Specifically, the employed annotation guidelines involve the instructions for using the annotation platform and detailed guidelines about the annotation procedure, and we demonstrate them in Table 9.
| User Guidelines |
| Part I Introduction Welcome to the annotation platform. This platform is designed to simplify the annotation process and enhance annotation efficiency. Before the detailed introduction, we want to claim again that you may feel inconvenient as the evaluated models may generate images with uncomfortable and inappropriate visual content. Now, you are still welcomed if you want to withdraw your consent The annotation process is conducted on a sample-by-sample basis, with a question-by-question approach. Thus, you are supposed to answer all the questions raised for the present sample to accomplish its annotation. Once all the delegated samples are accomplished, your job is finished and we are thankful for your contribution to the project. |
| Part II Guidelines of the User Interface 1.User Login: To access the annotation platform, you are required to login as a user. Please navigate to the login page, enter the username and password provided by us, and click the “Login” button. 2.Dashboard: Once you complete the login, you will be jumped into the dashborad page. The dashboard will list the overview of the samples assigned to you to annotate. Besides, we list the status of each sample for you to freely check your annotation progress (e.g., pending, completed). 3.Annotation Interface: Click on the “Start” button or an assigned image through the dashboard interface, you will jump into the annotation interface. annotation interface is made up of three components: 1) Image Display: View the image to be annotated and its conditioned prompts; 2) Question Panel: List of single-choice questions related to the image; 3) Navigation Buttons: "Next" and "Previous" buttons to navigate through questions and images. 4.Answering Questions: Each time, the annotation interface will provide you a sample for annotation, please view the image and read the associated question, select the appropriate answer from the available options, and repeat the processs for all questions related to the question. 5.Saving and Submitting Annotations: To save progress and submit completed annotations, you can click the “Save” button to save your progress. If you finish the assigned sample and ensure the accuracy and confidence of its, you can click the “Submit” button to submit this annotation. 6.Review and Edit Annotations: If you want to review and edit your submission, you can navigate the completed tasks section, and select the image to review. You will jump into its annotation interface with the previously submitted annotations and are allowed to do any modification. 7.Report and contact: If you find any problem about the assigned sample, such as witnessing NSFW or biased content, assigned visually abnormal sample, feel free to click the “Report” button and fill a form to report this sample. If you have any question about the standard of the annotation or have suggestions for improvement, please do not hesitate to contact us through phone, we will be glad to help you. |
| Part III General Guidelines of the Human Annotation 1.In general, you are supposed to answer all the questions raised for the present sample to accomplish its annotation. This annotation only involves single-choice question. 2.Before answering the question, please ensure that the question is applicable to this prompt. If it is not applicable, please select option 0 directly—this is the predefined option for this particular scenario. 3.If you are answering a question about the image faithfulness, you may find the question is applicable to multiple objects within the image. you need to answer the question regarding to every applicable object and its role in the image. A straightforward way for this is to solely score every applicable object and choose the option closest to the calculated weighted average score. 4.If you are answering the object faithfulness question on the image faithfulness annotation, you need to drop and report for the encountered image with no clear main object. 5.If you are answering the commonsense question on the image faithfulness annotation, you need to drop and report for the encountered surreal and sci-fi image. 6.You are required to first annotate 30 samples to form a stable and reasonable assessment standard. Then, accomplish the annotation in progress. 7.This annotation is for evaluating image faithfulness and text-image alignment, as a consequence, the standard of the annotation is universal. 8.If you feel confused at anything about the human annotation, feel free to contact us through phone, we will be glad to help you. 9. Once you have submitted your annotation results, we are very thankful to inform you that you have finished your job. Thank you once again for your contribution to our project. 10.If you have submitted your annotation but want to withdraw your submission and review the annotation results, you can contact us through phone, we will send it back to you. |
Trial Annotation Even with the above preparation, there is no quantitative evidence to verify the quality, the efficiency, and the inter-annotator consistency of the human annotation. Additionally, the standard for assessing image faithfulness and text-image are universal, which further emphasize the significant role of high inter-annotator consistency. Considering that, we conduct a multi-turn trial annotation on another 50 synthesized images. After each trial, we calculate the Cohen’s kappa coefficient and conduct a meeting for our annotators to explain annotation standards, rules and guidelines, thereby ensuring high inter-annotator reliability. In total, we conduct nine turns of trial annotation, and in the last turn of the trial, the Cohen’s kappa coefficient of our annotators reaches , indicating high inter-annotator reliability.
Random Sampling Quality Inspection Upon reaching the milestone percentages of 25%, 50%, 75%, and 100% in the annotation progress, we conducted a series of random sampling quality inspections on the present annotation results at each milestone, totally four turns of random sampling quality inspection. The random sampling quality inspection by four experts in text-to-image generation selected from our group on 1,000 randomly sampled annotated images. For the first two turn of quality inspection, there are totally 423 and 112 annotated samples that failed the inspection. The failed samples are re-annotated and re-inspected. As for the last two turns of quality inspection, they both revealed zero failed samples due to the thoughtful and rigorous annotation preparation.
Appendix D Additional Details of the Evaluated Models
In this section, we introduce the details of the evaluated text-to-image generative models in this work.
Stable Diffusion {v1.4, v1.5, v2 base, v2.0, v2.1}. Stable Diffusion (SD) is a series of 1B text-to-image generative models based on latent diffusion model [41] and is trained on LAION-5B [45]. Specifically, the SD series includes SD v1.1, SD v1.2, SD v1.4, SD v1.5, SD v2 base, SD v2.0, and SD v2.1 respectively. Among them, we choose the most commonly-employed SD v1.4, SD v1.5, SD v2.0 and SD v2.1 for EvalAlign evaluation.
SD v1.1 was trained at a resolution of 256x256 on laion2B-en for 237k steps, followed by training at a resolution of 512x512 on laion-high-resolution ((170M examples from LAION-5B with resolution >= 1024x1024) for the subsequent 194k steps. While, SD v1.2 was initialized from v1.1 and further finetuned for 515k steps at resolution 512x512 on laion-aesthetics v2 5+ (a subset of laion2B-en, filtered to images with an original size >= 512x512, estimated aesthetics score > 5.0, and an estimated watermark probability < 0.5). SD v1.4 is initialized from v1.2 and subsequently finetuned for 225k steps at resolution 512x512 on laion-aesthetics v2 5+. This version incorporates a 10% dropping of the text-conditioning to improve classifier-free guidance sampling. Similar to SD v1.4, SD v1.5 is resumed from SD v1.2 and trained 595k steps at resolution 512x512 on laion-aesthetics v2 5+, with 10% dropping of the text-conditioning.
SD v2 base is trained from scratch for 550k steps at resolution 256x256 on a subset of LAION-5B filtered for explicit pornographic material, using the LAION-NSFW classifier with punsafe = 0.1 and an aesthetic score >= 4.5. Then it is further trained for 850k steps at resolution 512x512 on the same dataset on images with resolution >= 512x512. SD v2.0 is resumed from stable-diffusion v2 base and trained for 150k steps using a v-objective on the same dataset. After that, it is further finetuned for another 140k steps on 768x768 images. SD v2.1 is finetuned from SD v2.0 with an additional 55k steps on the same dataset (with punsafe=0.1), and then finetuned for another 155k extra steps with punsafe=0.98.
Stable Diffusion XL {v1.0, Refiner v1.0}. Stable Diffusion XL (SDXL) is a powerful text-to-image generation model that iterates on the previous Stable Diffusion models in three key ways: (1) its UNet is 3x larger and SDXL combines a second text encoder (OpenCLIP ViT-bigG/14) with the original text encoder to significantly increase the number of parameters; (2) it introduces size and crop-conditioning to preserve training data from being discarded and gain more control over how a generated image should be cropped; (3) it introduces a two-stage model process; the base model (can also be run as a standalone model) generates an image as an input to the refiner model which adds additional high-quality details.
Pixart-Alpha. Pixart-Alpha is a model that can be used to generate and modify images based on text prompts. It is a Transformer Latent Diffusion Model that uses one fixed, pretrained text encoders (T5)) and one latent feature encoder (VAE).
Latent Consistency Model Stable Diffusion XL Latent Consistency Model Stable Diffusion XL (LCM SDXL) [28] enables SDXL for swift inference with minimal steps. Viewing the guided reverse diffusion process as solving an augmented probability flow ODE (PF-ODE), LCMs are designed to directly predict the solution of such ODE in latent space, mitigating the need for numerous iterations and allowing rapid, high-fidelity sampling.
Dreamlike Diffusion 1.0. Dreamlike Diffusion 1.0 [10] is a SD v1.5 model finetuned on high-quality art images by dreamlike.art.
Dreamlike Photoreal 2.0. Dreamlike Photoreal 2.0 [11] is a photorealistic text-to-image latent diffusion model resumed from SD v1.5 by dreamlike art. This model was finetuned on 768x768 images, it works pretty good with resolution 768x768, 640x896, 896x640 and higher resolution such as 768x1024.
Redshift Diffusion. Redshift Diffusion [39] is a Stable Diffusion model finetuned on high-resolution 3D artworks.
Vintedois Diffusion. Vintedois Diffusion [51] is a Stable Diffusion v1.5 model finetuned on a large number of high-quality images with simple prompts to generate beautiful images without a lot of prompt engineering.
Safe Stable Diffusion {Weak, Medium, Strong, Max}. Safe Stable Diffusion [31] is an enhanced version of the SD v1.5 model by mitigating inappropriate degeneration caused by pretraining on unfiltered web-crawled datasets. For instance SD may unexpectedly generate nudity, violence, images depicting self-harm, and otherwise offensive content. Safe Stable Diffusion is an extension of Stable Diffusion that drastically reduces this type of content. Specifically, it has an additional safety guidance mechanism that aims to suppress and remove inappropriate content (hate, harassment, violence, self-harm, sexual content, shocking images, and illegal activity) during image generation. The strength levels for inappropriate content removal are categorized as: {Weak, Medium, Strong, Max}.
MultiFusion. MultiFusion [29] is a multimodal, multilingual diffusion model that extends the capabilities of SD v1.4 by integrating various modules to transfer capabilities to the downstream model. This combination results in novel decoder embeddings, which enable prompting of the image generation model with interleaved multimodal, multilingual inputs, despite being trained solely on monomodal data in a single language.
DeepFloyd-IF { M, L, XL } v1.0. DeepFloyd-IF [1] is a novel state-of-the-art open-source text-to-image model with a high degree of photorealism and language understanding. It is a modular composed of a frozen text encoder and three cascaded pixel diffusion modules: a base model that generates 64x64 image based on text prompt and two super-resolution models, each designed to generate images of increasing resolution: 256x256 and 1024x1024, respectively. All stages of the model utilize a frozen text encoder based on the T5 transformer to extract text embeddings, which are then fed into a UNet architecture enhanced with cross-attention and attention pooling. Besides, it underscores the potential of larger UNet architectures in the first stage of cascaded diffusion models and depicts a promising future for text-to-image synthesis. The model is available in three different sizes: M, L, and XL. M has 0.4B parameters, L has 0.9B parameters, and XL has 4.3B parameters.
Throughout all of the conducted experiments, we use the default inference settings for each evaluated models.
Appendix E Instruction Templates
Here, we present every instruction used for EvalAlign evaluation on image faithfuleness and text-image alignment. The templates contain some placeholders set for filling in the corresponding attributes of the input images during the evaluation. For example, a specific “<ObjectHere>” and “<NumberHere>” can be “people, laptop, scissors.” and “plate: 1, turkey sandwich: 3, lettuce: 1.”, respectively.
For EvalAlign evaluation on image faithfulness, we devise 5 questions concentrate on the faithfulness of the generated body structure, generated face, generated hand, generated objects, as well as generation adherence to commonsense and logic. The instruction templates for these fine-grained criteria are as follows:
The templates of EvalAlign evaluation on text-image alignment are as follows. We select 6 common aspects of text-image alignment, object, number, color, style, spatial relationship and action. For images that do not involve the specified attribute, the corresponding question template is not filled in and subsequently input into EvalAlign.
Appendix F Additional Quantitative Analysis
F.1 Generalization Experiments
To verify the generalization capability of our evaluation model, We compared MLLM’s SFT using different training datasets: one with images generated by all 8 text-to-image models and another with images generated by only 4 of these models, while the final evaluation was conducted on images generated by the other 4 models. As shown in Table 10 and Table 11, We observed that MLLMs trained on images from a subset of text-to-image models can effectively generalize to images generated by unseen text-to-image models.
| Method | T2I models | body | hand | face | object | common | MAE |
| Human | - | 1.4988 | 0.8638 | 1.1648 | 2.2096 | 0.8710 | 0 |
| EvalAlign | 8 | 1.6058 | 0.7901 | 1.1974 | 2.2783 | 0.8871 | 0.0596 |
| 4 | 1.6522 | 0.9588 | 1.2355 | 2.3032 | 0.9516 | 0.0987 |
| Method | T2I models | Object | Count | Color | Style | Spatial | Action | MAE |
| Human | - | 1.7373 | 1.3131 | 2.0000 | 1.9333 | 1.5952 | 1.8837 | 0 |
| EvalAlign | 8 | 1.7203 | 1.3232 | 1.9565 | 1.9333 | 1.6547 | 1.8605 | 0.0256 |
| 4 | 1.7832 | 1.3526 | 1.9637 | 1.9876 | 1.6891 | 1.8954 | 0.0469 |
F.2 Instruction Enhancement Experiments
| Method | Instruction | SDXL | Pixart | Wuerstchen | SDXL-Turbo | IF | SD v1.5 | SD v2.1 | LCM | MAE |
| Human | – | 2.1044 | 1.8606 | 1.7839 | 1.3854 | 1.3822 | 1.3818 | 1.1766 | 1.0066 | 0 |
| EvalAlign | ✗ | 1.9565 | 1.9286 | 1.8565 | 1.1818 | 1.3419 | 1.4801 | 1.4078 | 1.1051 | 0.1201 |
| ✓ | 2.0443 | 1.9199 | 1.8012 | 1.3353 | 1.2960 | 1.4702 | 1.3221 | 1.0305 | 0.0064 |
Providing more contextual information for instructions enhances the performance of MLLMs. To further improve MLLM evaluation performance, we enhanced the prompts for both SFT and inference stages. As shown in Table 12, our experiments demonstrate that the enhanced prompts significantly increase evaluation accuracy. Specifically, the evaluation using enhanced instructions reduced the MAE metric by half, from 0.120 to 0.006, compared to the original instructions. Additionally, this approach consistently improved evaluation performance across different text-to-image models.
F.3 Mulit-scaling Resolutions Experiments
| Method | Multi Scale | SDXL | Pixart | Wuerstchen | SDXL-Turbo | IF | SD v1.5 | SD v2.1 | LCM | MAE |
| Human | – | 2.1044 | 1.8606 | 1.7839 | 1.3854 | 1.3822 | 1.3818 | 1.1766 | 1.0066 | 0 |
| EvalAlign | 1.8105 | 1.9238 | 1.9325 | 1.2078 | 1.2247 | 1.4540 | 1.3012 | 1.0554 | 0.1358 | |
| ✓ | 2.0443 | 1.9199 | 1.8012 | 1.3353 | 1.296 | 1.4702 | 1.3221 | 1.0305 | 0.0064 |
In the design of LLaVA-Next, using multi-scale resolution images as input helps address the issue of detail information loss, which significantly impacts the evaluation of image faithfulness, such as assessing deformations in hands and faces. We conducted a multi-scale image training comparison experiment to validate this approach. The baseline was the 13B LLaVA model with 336336 resolution input, while the comparison model used images at three resolutions (336336, 672672, 10081008) as input. As shown in Table 13, training with multi-scale inputs significantly enhanced the model’s understanding of image and achieved better evaluation performance.
F.4 Full Comparison with Existing Evaluation Methods
| Model | Human | EvalAlign | HPS v2 | CLIP-score | ImageReward | PickScore | IS |
| PixArt XL2 1024 MS [6] | 2.2848 | 1.6415 | 31.6226 | 0.8580 | 0.9696 | 22.1335 | 5.258321 |
| Dreamlike Photoreal v2.0 [11] | 2.0070 | 1.4522 | 29.2322 | 0.828612 | 0.1886 13 | 21.2271 | 5.645214 |
| SDXL Refiner v1.0 [47] | 1.9229 | 1.6072 | 29.8197 | 0.8566 | 0.7245 | 22.0492 | 5.398516 |
| SDXL v1.0 [33] | 1.8136 | 1.4675 | 29.0620 | 0.8467 | 0.7043 | 21.8106 | 5.377418 |
| Wuerstchen [32] | 1.7837 | 1.4279 | 30.6622 | 0.819914 | 0.3212 11 | 21.3720 | 5.9751 |
| LCM SDXL [28] | 1.6910 | 1.3391 | 29.3588 | 0.833510 | 0.5304 | 21.6532 | 5.9874 |
| Openjourney [34] | 1.6667 | 1.175010 | 26.347513 | 0.819615 | 0.1478 16 | 20.863710 | 5.736810 |
| Safe SD MAX [31] | 1.6491 | 1.2175 | 25.739617 | 0.755524 | -0.050722 | 20.459421 | 5.542815 |
| LCM LORA SDXL [28] | 1.6387 | 1.3833 | 27.329910 | 0.8364 | 0.4959 | 21.4824 | 5.657512 |
| Safe SD STRONG [31] | 1.630810 | 1.146611 | 25.576418 | 0.816518 | -0.102223 | 20.621118 | 5.8643 |
| Safe SD MEDIUM [31] | 1.627511 | 1.129815 | 26.279814 | 0.810120 | 0.2042 12 | 20.788012 | 6.1760 |
| Safe SD WEAK [31] | 1.607812 | 1.118817 | 26.118015 | 0.780923 | -0.126424 | 20.387324 | 5.386117 |
| SD v2.1 [40] | 1.552413 | 1.109418 | 26.582312 | 0.8377 | 0.4116 | 21.0502 | 5.307319 |
| SD v2.0 [40] | 1.527714 | 1.130014 | 25.348121 | 0.817017 | 0.0872 18 | 20.752913 | 5.9060 |
| Openjourney v2 [35] | 1.500015 | 0.995620 | 24.698423 | 0.795822 | -0.041521 | 20.408822 | 5.199522 |
| Redshift diffusion [39] | 1.473316 | 1.138212 | 25.157222 | 0.810121 | 0.0218 20 | 20.615519 | 5.7657 |
| Dreamlike Diffusion v1.0 [10] | 1.465217 | 1.2052 | 29.6506 | 0.8543 | 0.6508 | 21.2664 | 5.661411 |
| SD v1.5 [40] | 1.441718 | 1.136213 | 25.497219 | 0.821413 | 0.1686 14 | 20.714316 | 6.0535 |
| IF-I-XL v1.0 [1] | 1.380819 | 0.922122 | 27.4512 | 0.8449 | 0.6087 | 20.747414 | 5.301220 |
| SD v1.4 [40] | 1.359220 | 0.951121 | 25.369720 | 0.819016 | 0.1050 17 | 20.653517 | 5.8571 |
| Vintedois Diffusion v0.1 [51] | 1.356221 | 1.079719 | 26.590111 | 0.8341 | 0.3562 10 | 20.835811 | 5.018123 |
| IF-I-L v1.0 [1] | 1.263522 | 0.881423 | 27.4836 | 0.8384 | 0.4463 | 20.717015 | 5.650213 |
| MultiFusion [29] | 1.237223 | 1.129816 | 23.813324 | 0.815119 | 0.0695 19 | 20.478020 | 4.382424 |
| IF-I-M v1.0 [1] | 1.013524 | 0.792824 | 25.952216 | 0.832911 | 0.1637 15 | 20.403523 | 5.7451 |
| Model | Human | EvalAlign | HPS v2 | CLIP-score | ImageReward | PickScore | IS |
| IF-I-XL v1.0 [1] | 5.4500 | 5.5300 | 32.547710 | 0.8579 | 0.4391 | 21.199810 | 7.1864 |
| IF-I-L v1.0 [1] | 5.2300 | 5.4500 | 32.7140 | 0.8538 | 0.3820 | 21.128412 | 6.657120 |
| SDXL Refiner v1.0 [47] | 5.2100 | 5.4000 | 35.6465 | 0.8528 | 0.4738 | 22.3532 | 7.1101 |
| LCM SDXL [28] | 5.1800 | 5.3300 | 33.8011 | 0.8512 | 0.3833 | 21.9620 | 7.006714 |
| PixArt XL2 1024 MS [6] | 5.1100 | 5.3100 | 37.0493 | 0.8634 | 0.6542 | 22.3926 | 6.929316 |
| IF-I-M v1.0 [1] | 5.0800 | 5.2200 | 31.095114 | 0.8434 | 0.0499 10 | 20.827020 | 7.2586 |
| LCM LORA SDXL [28] | 5.0600 | 5.2700 | 32.7752 | 0.834910 | 0.1618 | 21.7627 | 6.838919 |
| SDXL v1.0 [33] | 5.0300 | 5.3500 | 35.1593 | 0.8540 | 0.4322 | 22.1291 | 7.2762 |
| Wuerstchen [32] | 4.8700 | 5.1700 | 36.4632 | 0.8381 | 0.2513 | 21.7779 | 7.1280 |
| Openjourney [34] | 4.830010 | 4.920015 | 31.149512 | 0.817316 | -0.086714 | 21.116313 | 6.272922 |
| SD v2.1 [40] | 4.800011 | 5.070011 | 31.101713 | 0.827814 | -0.045312 | 21.2093 | 7.0767 |
| MultiFusion [29] | 4.680012 | 4.800018 | 28.795724 | 0.826415 | -0.133715 | 20.962517 | 4.959324 |
| Dreamlike Diffusion v1.0 [10] | 4.660013 | 5.150010 | 34.8196 | 0.8493 | 0.2295 | 21.5550 | 6.858118 |
| SD v2.0 [40] | 4.640014 | 5.010012 | 30.615317 | 0.829813 | -0.142416 | 21.190511 | 7.012413 |
| Vintedois Diffusion v0.1 [51] | 4.620015 | 4.950014 | 31.950311 | 0.831912 | -0.022211 | 21.114114 | 6.198723 |
| Safe SD STRONG [31] | 4.600016 | 4.830017 | 30.661516 | 0.775123 | -0.502822 | 20.749121 | 7.2743 |
| Dreamlike Photoreal v2.0 [11] | 4.560017 | 4.980013 | 33.7712 | 0.834411 | -0.085913 | 21.4832 | 6.936315 |
| Safe SD WEAK [31] | 4.530018 | 4.710020 | 30.564418 | 0.814018 | -0.272818 | 20.989916 | 6.423821 |
| SD v1.4 [40] | 4.520019 | 4.760019 | 29.914920 | 0.804820 | -0.343819 | 20.846219 | 7.015012 |
| SD v1.5 [40] | 4.450020 | 4.900016 | 30.167319 | 0.814217 | -0.221317 | 20.864018 | 7.1642 |
| Safe SD MEDIUM [31] | 4.400021 | 4.560024 | 30.782015 | 0.797421 | -0.359120 | 21.025715 | 7.070910 |
| Redshift diffusion [39] | 4.350022 | 4.670021 | 29.286522 | 0.806619 | -0.417221 | 20.632723 | 7.061011 |
| Safe SD MAX [31] | 4.310023 | 4.590023 | 29.812621 | 0.760124 | -0.609524 | 20.704622 | 7.2273 |
| Openjourney v2 [35] | 4.150024 | 4.650022 | 29.238923 | 0.785122 | -0.605123 | 20.597324 | 6.861317 |
Table 14 and Table 15 present the comparison results between our proposed method and existing alternatives. Owing to the powerful aligned understanding of images and text by MLLMs, our model achieves the best alignment with human performance across all tested benchmarks for 24 text-to-image models, along with a significant improvement in image faithfulness. This indicates that our model excels in both comprehending and evaluating the intricate details of generated images, closely mirroring human judgment and setting a new standard for image-text alignment and faithfulness.
Appendix G Qualitative Analysis
As shown in Figure 5, we present a comparison of different evaluation metrics on images generated by four models, including human annotated scores, EvalAlign, ImageReward [56], HPSv2 [54], and PickScore [20]. The digits in the figure represent the ranking for each evaluation metric, with darker colors indicating higher rankings. From the figure, it is evident that our proposed EvalAlign metric closely matches the human rankings across two evaluation dimensions, demonstrating excellent consistency.
Furthermore, Figure 6 showcases the EvalAlign evaluation metric across different fine-grained results. The numbers in the figure represent EvalAlign scores for the corresponding evaluation aspect, with darker colors indicating higher scores and better generation performance. Note that if the text prompt does not specify a particular style, the style consistency score defaults to 0. From these results, it is evident that the same text-to-image model exhibits significant performance variation across different evaluation aspects.
Appendix H Datasheet
H.1 Motivation
Q1
For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
Existing benchmarks for text-to-image generation models are neither comprehensive nor accurate enough to meet current evaluation needs. The reasons for this include: (1) Limited model parameters: Current evaluation models have too few parameters, which restricts their ability to accurately represent images, leading to significant discrepancies compared to human evaluations. (2) Training data limitations: Some evaluation methods, such as Inception Score (IS) [44], Frechet Inception Distance (FID) [15], and CLIP Score [14], use models that have not been trained with synthesized images, which may introduce training bias and flaws the evaluation. (3) High annotation costs: Some methods, such as ImageReward [56], HPS [55] and HPS v2 [54], rely heavily on extensive human annotations, which significantly increases the cost of labeling. (4) Lack of detailed evaluation metric: The evaluation metrics do not provide fine-grained interpretability, preventing them from guiding model optimization effectively. (5) Computational inefficiency: The evaluation models require substantial computational resources, making them inefficient. The EvalAlign benchmark fills in this gap, as it offers low-cost, accurate, and efficient model evaluations while providing fine-grained, interpretable metrics.
Q2
Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
This benchmark is presented by Fudan University, we concentrate on AI for meteorological, image and video generation, multimodal large language model. Fudan-FUXI is affiliated with Fudan University and led by Prof. Hao Li.
Q3
Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
This work was supported by Fudan University.
Q4
Any other comments?
No.
H.2 Composition
Q5
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
EvalAlign benchmark provides two sets of prompts. The evaluation is based on fine-grained scoring of the images generated by the evaluated models and conditioned on the provided prompts. These two sets of prompts are used for the EvalAlign evaluation on image faithfulness and text-image alignment, respectively.
Q6
How many instances are there in total (of each type, if appropriate)?
We provide 1,500 prompts meticulously selected for the EvalAlign evaluation on image faithfulness, and another 1,500 prompts for the EvalAlign evaluation on text-image alignment.
Q7
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
Yes. The aforementioned prompts are all sourced from the utilized prompts are soureced from HPS [55], HRS-Bench [2], HPSv2 [54], TIFA [16], DSG [7], T2I-Comp [17], Winoground [50], DALL-EVAL [8], DiffusionDB [52], PartiPrompts [57], DrawBench [43], and JourneryDB [49]. we filter the prompts from these sources to ensure the quantity, quality and diversity of the two cultivated prompt sets. For the image faithfulness, we prioritize prompts related to human, animals, and other tangible objects, as prompts depicting sci-fi scenarios are less suitable for this type of assessment. The selected prompts encompass 10k subjects across 15 categories. For the text-image alignment, we refined our selection based on descriptions of style, color, quantity, and spatial relationships in the prompts. Only those prompts containing relevant descriptions and exceeding 15 words in length were considered.
Q8
What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
The instances of the two prompt sets are typical textual prompts used as guidance for text-to-image generation.
Q9
Is there a label or target associated with each instance? If so, please provide a description.
No, the prompts are utilized as guidance for text-to-image generation. EvalAlign can score the synthesized images conditioned on these prompts without the need of reference images.
Q10
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
No. There is nothing missed from individual instances.
Q11
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
Yes. There are totally 3,000 prompts, half of them is used for EvalAlign evaluation on image faithfulness, while the other half is used for EvalAlign evaluation on text-image alignment.
Q12
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
No.
Q13
Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
No.
Q14
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
The dataset is self-contained. Everything is available at https://github.com/SAIS-FUXI/EvalAlign.
Q15
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)? If so, please provide a description.
No. All of the released prompts are sourced from existing open-source datasets and filtered by our prompt filtering pipelines. The sourced datasets have already cleaned any possible confidential. Besides, our prompt filtering pipeline also includes a confidential cleaning stage.
Q16
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
No. All of the released prompts are filtered by our prompt filtering pipelines, which includes a NSFW filtering stage to filter out prompts that might be offensive, insulting, threatening, or might otherwise cause anxiety.
Q17
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
People may be present in the prompts, but people are not the sole focus of the dataset.
Q18
Does the dataset identify any subpopulations (e.g., by age, gender)?
We use automated gender and skin tone classifiers to identify the filtered prompts for these information. After that, the filtered prompts are randomly sampled to erase these biases.
Q19
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
No. We only provide textual prompts, and we have cleaned the possible confidential information contained in these prompts.
Q20
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)? If so, please provide a description.
No. They are considered as private content and have been cleaned from our datasets.
Q21
Any other comments?
We caution discretion on behalf of the user and call for responsible usage of the benchmark for research purposes only.
H.3 Collection Process
Q22
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
The released data is observable raw text.
Q23
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
The prompts are collected from popular open-source datasets.
Q24
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
The prompts are first sampled to erase possible biases such as gender, skin tone and region, and are then randomly sampled to 1,500 and 1,500 for the EvalAlign evaluation on image faithfulness and text-image alignment.
Q25
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
The authors of this paper downloaded the released prompt sets.
Crowdworkers were only involved for annotating instruction tuning data for EvalAlign, which is not prepared for release.
Q26
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
The data was collected from January 2024 to June 2024.
Q27
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
Yes, we conduct data filtering by human annotators.
Q28
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
People may appear in the prompts but they are not the only focus of the dataset.
Q29
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
Our prompts were collected from third party websites.
Q30
Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
Individuals involved in SFT training data annotation were notified about the data collection. We used a specially designed annotation user interface, and conducted annotation training and trial annotation to explain every aspect about the data annotation.
Q31
Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
Yes.
Q32
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
Yes.
Q33
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
Yes, we discuss the limitation of EvalAlign in appendix A.
Q34
Any other comments?
No.
H.4 Preprocessing, Cleaning, and/or Labeling
Q35
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remainder of the questions in this section.
Yes, we filter collected prompts with biased content, NSFW content, and low quality.
Q36
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
Yes, the “raw” data can be obtained form the release repository of the sourced prompt datasets.
Q37
Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
Yes. Our prompt cleaning pipeline is constructed on open source Github projects.
Q38
Any other comments?
No.
H.5 Uses
Q39
Has the dataset been used for any tasks already? If so, please provide a description.
No. EvalAlign is a new benchmark for text-to-image generation.
Q40
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
Yes. The data, model weight and code utilized by the paper will be available at https://github.com/SAIS-FUXI/EvalAlign.
Q41
What (other) tasks could the dataset be used for?
The main usage of our benchmark is text-to-image generation.
Q42
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description. Is there anything a future user could do to mitigate these undesirable harms?
The released data has removed bias such as gender and skin tone.
However, with the development of society and change of human cognition, we can’t ensure that the biases recognised in the future can be removed.
We emphasize that this dataset is permitted for use solely for research purposes.
Q43
Are there tasks for which the dataset should not be used? If so, please provide a description.
Because the model-generated images conditioned on the provided prompts may contain bias and toxic content, under no circumstance should these images or models trained on them be put into production. It is neither safe nor responsible.
The dataset should be solely used for research purposes.
Q44
Any other comments?
No.
H.6 Distribution and License
Q45
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
Yes, EvalAlign will be open-source.
Q46
How will the dataset be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
Our code, model weight and data will be released at https://github.com/SAIS-FUXI/EvalAlign.
Q47
When will the dataset be distributed?
June 13, 2024 and onward.
Q48
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
The released prompts are sourced from existing open-source datasets. The license associated with them is followed accordingly.
We also release the evaluation model weights and the code used in the paper under the Apache-2.0 license.
Q49
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
We do not own the copyright of the released prompts.
Q50
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
No.
Q51
Any other comments?
No.
H.7 Maintenance
Q52
Who will be supporting/hosting/maintaining the dataset?
Fudan University, the builder and owner of the EvalAlign benchmark, will be supporting, hosting, and maintaining the project.
Q53
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Pleaser refer to http://www.fudan.edu.cn
Q54
Is there an erratum? If so, please provide a link or other access point.
No.
Q55
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
Yes, the progress of this will be documented at https://github.com/SAIS-FUXI/EvalAlign.
Q56
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
Yes, People may contact us to add specific samples to a blacklist.
Q57
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
Yes, we will archive the older version for the community to use and host other versions.
Q58
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to other users? If so, please provide a description.
People may contact us at fudan.edu.cn to request adding new prompts, models and codes.
Q59
Any other comments?
No.

### v2

@@
 | EvalAlign | 1.6807 | 1.2516 | 1.8696 | 1.9592 | 1.5882 | 1.8382 |
+| Method | Data Size | SDXL | Pixart | Wuerstchen | SDXL-Turbo | IF | SD v1.5 | SD v2.1 | LCM |
+| Human | – | 2.1044 | 1.8606 | 1.7839 | 1.3854 | 1.3822 | 1.3818 | 1.1766 | 1.0066 |
+| EvalAlign | 200 | 1.7443 | 1.8898 | 1.9278 | 1.1261 | 1.2977 | 1.5254 | 1.4309 | 1.1204 |
+| 500 | 1.8890 | 1.9161 | 1.8586 | 1.2141 | 1.3109 | 1.3926 | 1.3815 | 0.9485 | |
+| 800 | 2.0443 | 1.9199 | 1.8012 | 1.3353 | 1.296 | 1.4702 | 1.3221 | 1.0305 |
+| Method | Model Size | SDXL | Pixart | Wuerstchen | SDXL-Turbo | IF | SD v1.5 | SD v2.1 | LCM |
+| Human | – | 2.1044 | 1.8606 | 1.7839 | 1.3854 | 1.3822 | 1.3818 | 1.1766 | 1.0066 |
+| EvalAlign | 7B | 1.9959 | 1.8615 | 1.8228 | 1.1708 | 1.2704 | 1.4031 | 1.3063 | 1.0145 |
+| 13B | 2.0443 | 1.9199 | 1.8012 | 1.3353 | 1.2960 | 1.4702 | 1.3221 | 1.0305 | |
+| 34B | 2.1131 | 1.8621 | 1.8083 | 1.3906 | 1.3076 | 1.3921 | 1.2037 | 1.0143 |
 Effect of dataset size for vision-language model training.
@@
 This result suggests that our method requires only a small amount of annotated data to achieve good performance, highlighting the cost-effectiveness of our approach.
-| Method | Data Size | SDXL | Pixart | Wuerstchen | SDXL-Turbo | IF | SD v1.5 | SD v2.1 | LCM |
-| Human | – | 2.1044 | 1.8606 | 1.7839 | 1.3854 | 1.3822 | 1.3818 | 1.1766 | 1.0066 |
-| EvalAlign | 200 | 1.7443 | 1.8898 | 1.9278 | 1.1261 | 1.2977 | 1.5254 | 1.4309 | 1.1204 |
-| 500 | 1.8890 | 1.9161 | 1.8586 | 1.2141 | 1.3109 | 1.3926 | 1.3815 | 0.9485 | |
-| 800 | 2.0443 | 1.9199 | 1.8012 | 1.3353 | 1.296 | 1.4702 | 1.3221 | 1.0305 |
 Effect of model size.
@@
 Specifically, increasing the MLLM model size from 7B to 34B results in notable improvements in performance.
-| Method | Model Size | SDXL | Pixart | Wuerstchen | SDXL-Turbo | IF | SD v1.5 | SD v2.1 | LCM |
-| Human | – | 2.1044 | 1.8606 | 1.7839 | 1.3854 | 1.3822 | 1.3818 | 1.1766 | 1.0066 |
-| EvalAlign | 7B | 1.9959 | 1.8615 | 1.8228 | 1.1708 | 1.2704 | 1.4031 | 1.3063 | 1.0145 |
-| 13B | 2.0443 | 1.9199 | 1.8012 | 1.3353 | 1.2960 | 1.4702 | 1.3221 | 1.0305 | |
-| 34B | 2.1131 | 1.8621 | 1.8083 | 1.3906 | 1.3076 | 1.3921 | 1.2037 | 1.0143 |
 5.3 Comparison with Existing Evaluation Methods
@@
 This experimental results confirms that our SFT training enables the MLLM to be successfully applied to the task of evaluating text-to-image models.
+Comparison with state-of-the-art methods.
+To verify the human preference alignment of our model, especially when compared with other baseline methods, we calculate Kendall rank [19] and Pearson [12] correlation coefficient and summarize the results in Table 8.
+This comparison experiment encompasses images generated by 24 text-to-image models, covering an extensive range of synthesized image distribution.
 | Method | Faithfulness | Alignment | ||
@@
 | EvalAlign | 0.7464 | 0.8730 | 0.8043 | 0.9356 |
-Comparison with state-of-the-art methods.
-To verify the human preference alignment of our model, especially when compared with other baseline methods, we calculate Kendall rank [19] and Pearson [12] correlation coefficient and summarize the results in Table 8.
-This comparison experiment encompasses images generated by 24 text-to-image models, covering an extensive range of synthesized image distribution.
 As can be concluded, compared with baseline methods, EvalAlign achieves significant higher alignment with fine-grained human preference on image faithfulness and image-text consistency, showcasing robust generalization ability.
@@
 Besides, since CLIP-s only cares the CLIP similarity of the generated image and its corresponding prompt, it behaves poorly in image faithfulness evaluation.
-The Per-question alignment and the final benchmark leaderboard of EvalAlign will be introduced in the supplementary materials.
+The per-question alignment and the leaderboard of EvalAlign will be introduced in the supplementary materials.
 6 Conclusion and Discussion
@@
 We hope that our work can inspire and facilitate future research in this field.
-Checklist
-The checklist follows the references.
-Please read the checklist guidelines carefully for information on how to answer these questions.
-For each question, change the default [TODO] to [Yes] , [No] , or [N/A] .
-You are strongly encouraged to include a justification to your answer, either by referencing the appropriate section of your paper or providing a brief inline description.
-For example:
-Did you include the license to the code and datasets? [Yes] See Section X.
-Did you include the license to the code and datasets? [No] The code and the data are proprietary.
-Did you include the license to the code and datasets? [N/A]
-Please do not modify the questions and only use the provided macros for your answers.
-Note that the Checklist section does not count towards the page limit.
-In your paper, please delete this instructions block and only keep the Checklist section heading above along with the questions/answers below.
-1.
-For all authors…
-Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
-Did you describe the limitations of your work? [Yes] The discussion of limitation is included in the supplementary files.
-Did you discuss any potential negative societal impacts of your work?
-[Yes] The discussion of negative societal impacts is recognised as part of the limitation, and will be included in the supplementary files.
-Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
-2.
-If you are including theoretical results…
-Did you state the full set of assumptions of all theoretical results? [N/A]
-Did you include complete proofs of all theoretical results? [N/A]
-3.
-If you ran experiments (e.g. for benchmarks)…
-Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
-[Yes] The instruction is given in the project page link of the paper.
-- (b)
-Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] All the experiments are reproduced by our members with different random seed.
-Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See 4.3.
-4.
-If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
-If your work uses existing assets, did you cite the creators? [Yes]
-Did you mention the license of the assets? [Yes] They will be mentioned in the Github repository if available.
-Did you include any new assets either in the supplemental material or as a URL? [Yes]
-Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes]
-Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See 3.2.
-5.
-If you used crowdsourcing or conducted research with human subjects…
-Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] They will be included in supplementary files.
-Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
-Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] The hourly wage for the works is 30 CNY/hour for every annotator.
 Appendix A Limitations
@@
 From these results, it is evident that the same text-to-image model exhibits significant performance variation across different evaluation aspects.
-Appendix H Datasheet
-H.1 Motivation
-Q1
-For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
-Existing benchmarks for text-to-image generation models are neither comprehensive nor accurate enough to meet current evaluation needs.
-The reasons for this include: (1) Limited model parameters: Current evaluation models have too few parameters, which restricts their ability to accurately represent images, leading to significant discrepancies compared to human evaluations.
-(2) Training data limitations: Some evaluation methods, such as Inception Score (IS) [44], Frechet Inception Distance (FID) [15], and CLIP Score [14], use models that have not been trained with synthesized images, which may introduce training bias and flaws the evaluation.
-(3) High annotation costs: Some methods, such as ImageReward [56], HPS [55] and HPS v2 [54], rely heavily on extensive human annotations, which significantly increases the cost of labeling.
-(4) Lack of detailed evaluation metric: The evaluation metrics do not provide fine-grained interpretability, preventing them from guiding model optimization effectively.
-(5) Computational inefficiency: The evaluation models require substantial computational resources, making them inefficient.
-The EvalAlign benchmark fills in this gap, as it offers low-cost, accurate, and efficient model evaluations while providing fine-grained, interpretable metrics.
-Q2
-Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
-This benchmark is presented by Fudan University, we concentrate on AI for meteorological, image and video generation, multimodal large language model.
-Fudan-FUXI is affiliated with Fudan University and led by Prof.
-Hao Li.
-Q3
-Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
-This work was supported by Fudan University.
-Q4
-Any other comments?
-No.
-H.2 Composition
-Q5
-What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)?
-Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)?
-Please provide a description.
-EvalAlign benchmark provides two sets of prompts.
-The evaluation is based on fine-grained scoring of the images generated by the evaluated models and conditioned on the provided prompts.
-These two sets of prompts are used for the EvalAlign evaluation on image faithfulness and text-image alignment, respectively.
-Q6
-How many instances are there in total (of each type, if appropriate)?
-We provide 1,500 prompts meticulously selected for the EvalAlign evaluation on image faithfulness, and another 1,500 prompts for the EvalAlign evaluation on text-image alignment.
-Q7
-Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
-If the dataset is a sample, then what is the larger set?
-Is the sample representative of the larger set (e.g., geographic coverage)?
-If so, please describe how this representativeness was validated/verified.
-If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
-Yes.
-The aforementioned prompts are all sourced from the utilized prompts are soureced from HPS [55], HRS-Bench [2], HPSv2 [54], TIFA [16], DSG [7], T2I-Comp [17], Winoground [50], DALL-EVAL [8], DiffusionDB [52], PartiPrompts [57], DrawBench [43], and JourneryDB [49].
-we filter the prompts from these sources to ensure the quantity, quality and diversity of the two cultivated prompt sets.
-For the image faithfulness, we prioritize prompts related to human, animals, and other tangible objects, as prompts depicting sci-fi scenarios are less suitable for this type of assessment.
-The selected prompts encompass 10k subjects across 15 categories.
-For the text-image alignment, we refined our selection based on descriptions of style, color, quantity, and spatial relationships in the prompts.
-Only those prompts containing relevant descriptions and exceeding 15 words in length were considered.
-Q8
-What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
-The instances of the two prompt sets are typical textual prompts used as guidance for text-to-image generation.
-Q9
-Is there a label or target associated with each instance? If so, please provide a description.
-No, the prompts are utilized as guidance for text-to-image generation. EvalAlign can score the synthesized images conditioned on these prompts without the need of reference images.
-Q10
-Is any information missing from individual instances?
-If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable).
-This does not include intentionally removed information, but might include, e.g., redacted text.
-No. There is nothing missed from individual instances.
-Q11
-Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
-Yes. There are totally 3,000 prompts, half of them is used for EvalAlign evaluation on image faithfulness, while the other half is used for EvalAlign evaluation on text-image alignment.
-Q12
-Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
-No.
-Q13
-Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
-No.
-Q14
-Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
-If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user?
-Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
-The dataset is self-contained. Everything is available at https://github.com/SAIS-FUXI/EvalAlign.
-Q15
-Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)?
-If so, please provide a description.
-No.
-All of the released prompts are sourced from existing open-source datasets and filtered by our prompt filtering pipelines.
-The sourced datasets have already cleaned any possible confidential.
-Besides, our prompt filtering pipeline also includes a confidential cleaning stage.
-Q16
-Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
-No.
-All of the released prompts are filtered by our prompt filtering pipelines, which includes a NSFW filtering stage to filter out prompts that might be offensive, insulting, threatening, or might otherwise cause anxiety.
-Q17
-Does the dataset relate to people? If not, you may skip the remaining questions in this section.
-People may be present in the prompts, but people are not the sole focus of the dataset.
-Q18
-Does the dataset identify any subpopulations (e.g., by age, gender)?
-We use automated gender and skin tone classifiers to identify the filtered prompts for these information. After that, the filtered prompts are randomly sampled to erase these biases.
-Q19
-Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
-No. We only provide textual prompts, and we have cleaned the possible confidential information contained in these prompts.
-Q20
-Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)?
-If so, please provide a description.
-No. They are considered as private content and have been cleaned from our datasets.
-Q21
-Any other comments?
-We caution discretion on behalf of the user and call for responsible usage of the benchmark for research purposes only.
-H.3 Collection Process
-Q22
-How was the data associated with each instance acquired?
-Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)?
-If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified?
-If so, please describe how.
-The released data is observable raw text.
-Q23
-What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)?
-How were these mechanisms or procedures validated?
-The prompts are collected from popular open-source datasets.
-Q24
-If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
-The prompts are first sampled to erase possible biases such as gender, skin tone and region, and are then randomly sampled to 1,500 and 1,500 for the EvalAlign evaluation on image faithfulness and text-image alignment.
-Q25
-Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
-The authors of this paper downloaded the released prompt sets.
-Crowdworkers were only involved for annotating instruction tuning data for EvalAlign, which is not prepared for release.
-Q26
-Over what timeframe was the data collected?
-Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)?
-If not, please describe the timeframe in which the data associated with the instances was created.
-The data was collected from January 2024 to June 2024.
-Q27
-Were any ethical review processes conducted (e.g., by an institutional review board)?
-If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
-Yes, we conduct data filtering by human annotators.
-Q28
-Does the dataset relate to people? If not, you may skip the remaining questions in this section.
-People may appear in the prompts but they are not the only focus of the dataset.
-Q29
-Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
-Our prompts were collected from third party websites.
-Q30
-Were the individuals in question notified about the data collection?
-If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
-Individuals involved in SFT training data annotation were notified about the data collection.
-We used a specially designed annotation user interface, and conducted annotation training and trial annotation to explain every aspect about the data annotation.
-Q31
-Did the individuals in question consent to the collection and use of their data?
-If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
-Yes.
-Q32
-If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses?
-If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
-Yes.
-Q33
-Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted?
-If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
-Yes, we discuss the limitation of EvalAlign in appendix A.
-Q34
-Any other comments?
-No.
-H.4 Preprocessing, Cleaning, and/or Labeling
-Q35
-Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?
-If so, please provide a description.
-If not, you may skip the remainder of the questions in this section.
-Yes, we filter collected prompts with biased content, NSFW content, and low quality.
-Q36
-Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
-Yes, the “raw” data can be obtained form the release repository of the sourced prompt datasets.
-Q37
-Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
-Yes. Our prompt cleaning pipeline is constructed on open source Github projects.
-Q38
-Any other comments?
-No.
-H.5 Uses
-Q39
-Has the dataset been used for any tasks already? If so, please provide a description.
-No. EvalAlign is a new benchmark for text-to-image generation.
-Q40
-Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
-Yes. The data, model weight and code utilized by the paper will be available at https://github.com/SAIS-FUXI/EvalAlign.
-Q41
-What (other) tasks could the dataset be used for?
-The main usage of our benchmark is text-to-image generation.
-Q42
-Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
-For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description.
-Is there anything a future user could do to mitigate these undesirable harms?
-The released data has removed bias such as gender and skin tone.
-However, with the development of society and change of human cognition, we can’t ensure that the biases recognised in the future can be removed.
-We emphasize that this dataset is permitted for use solely for research purposes.
-Q43
-Are there tasks for which the dataset should not be used? If so, please provide a description.
-Because the model-generated images conditioned on the provided prompts may contain bias and toxic content, under no circumstance should these images or models trained on them be put into production.
-It is neither safe nor responsible.
-The dataset should be solely used for research purposes.
-Q44
-Any other comments?
-No.
-H.6 Distribution and License
-Q45
-Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
-Yes, EvalAlign will be open-source.
-Q46
-How will the dataset be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
-Our code, model weight and data will be released at https://github.com/SAIS-FUXI/EvalAlign.
-Q47
-When will the dataset be distributed?
-June 13, 2024 and onward.
-Q48
-Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
-If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
-The released prompts are sourced from existing open-source datasets. The license associated with them is followed accordingly.
-We also release the evaluation model weights and the code used in the paper under the Apache-2.0 license.
-Q49
-Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
-If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
-We do not own the copyright of the released prompts.
-Q50
-Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
-If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
-No.
-Q51
-Any other comments?
-No.
-H.7 Maintenance
-Q52
-Who will be supporting/hosting/maintaining the dataset?
-Fudan University, the builder and owner of the EvalAlign benchmark, will be supporting, hosting, and maintaining the project.
-Q53
-How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
-Pleaser refer to http://www.fudan.edu.cn
-Q54
-Is there an erratum? If so, please provide a link or other access point.
-No.
-Q55
-Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
-If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
-Yes, the progress of this will be documented at https://github.com/SAIS-FUXI/EvalAlign.
-Q56
-If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
-If so, please describe these limits and explain how they will be enforced.
-Yes, People may contact us to add specific samples to a blacklist.
-Q57
-Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
-Yes, we will archive the older version for the community to use and host other versions.
-Q58
-If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
-If so, please provide a description.
-Will these contributions be validated/verified?
-If so, please describe how.
-If not, why not?
-Is there a process for communicating/distributing these contributions to other users?
-If so, please provide a description.
-People may contact us at fudan.edu.cn to request adding new prompts, models and codes.
-Q59
-Any other comments?
-No.

### v3

@@
 In this paper, we propose EvalAlign, a metric characterized by its accuracy, stability, and fine granularity.
-Our approach leverages the capabilities of Multimodal Large Language Models (MLLMs) pre-trained on extensive datasets.
+Our approach leverages the capabilities of Multimodal Large Language Models (MLLMs) pre-trained on extensive data.
 We develop evaluation protocols that focus on two key dimensions: image faithfulness and text-image alignment.
 Each protocol comprises a set of detailed, fine-grained instructions linked to specific scoring options, enabling precise manual scoring of the generated images.
-We Supervised Fine-Tune (SFT) the MLLM to align closely with human evaluative judgments, resulting in a robust evaluation model.
-Our comprehensive tests across 24 text-to-image generation models demonstrate that EvalAlign not only provides superior metric stability but also aligns more closely with human preferences than existing metrics, confirming its effectiveness and utility in model assessment.
+We supervised fine-tune (SFT) the MLLM to align with human evaluative judgments, resulting in a robust evaluation model.
+Our evaluation across 24 text-to-image generation models demonstrate that EvalAlign not only provides superior metric stability but also aligns more closely with human preferences than existing metrics, confirming its effectiveness and utility in model assessment.
 1 Introduction
-Text-to-image models, such as DALL·E series [38, 37, 3], Imagen [43], and Stable Diffusion [33], have significantly impacted various domains such as entertainment, design, and education, by enabling high-quality image generation.
-These technologies not only advance the field of text-to-image generation but also enhance related applications such as video generation [4, 60], image editing [46, 18, 59], and personalized image generation [13, 42].
-Despite this prevalence, in the practical use of generative models, we still lack evaluation methods to accurately evaluate existing text-to-image models.
-Existing benchmarks for text-to-image generation models are neither comprehensive nor accurate enough to meet current evaluation needs.
-The reasons for this include: (1) Limited model parameters: Current evaluation models have too few parameters, which restricts their ability to accurately represent images, leading to significant discrepancies compared to human evaluations.
-(2) Training data limitations: Some evaluation methods, such as Inception Score (IS) [44], Frechet Inception Distance (FID) [15], and CLIP Score [14], use models that have not been trained with synthesized images, which may introduce training bias and flaws the evaluation.
-(3) High annotation costs: Some methods, such as ImageReward [56], HPS [55] and HPS v2 [54], rely heavily on extensive human annotations, which significantly increases the cost of labeling.
-(4) Lack of detailed evaluation metric: The evaluation metrics do not provide fine-grained interpretability, preventing them from guiding model optimization effectively.
-(5) Computational inefficiency: The evaluation models require substantial computational resources, making them inefficient.
-To address these issues, we propose EvalAlign, which offers low-cost, accurate, and efficient model evaluations while providing fine-grained, interpretable metrics.
-Specifically, we Supervised Fine-Tune (SFT) a Multimodal Large Language Model (MLLM) to align it with human annotations.
-Our focus is on two key evaluation aspects: image faithfulness and text-image alignment.
-Due to its pre-training on large-scale datasets and a large number of model parameters, the MLLM demonstrates excellent understanding and generalization capabilities for images and instructions.
-This makes it possible to design evaluation instruction based on MLLM to evaluate text-to-image models.
-However, since the pre-trained datasets do not include model-generated images (such as distorted body structures and human hands) or evaluation-related text instructions, using MLLM directly for model evaluation does not yield optimal results.
-Therefore, we employ SFT on a small amount of manually annotated data to align MLLM with human annotations for text-to-image generation.
+Text-to-image models, such as DALL·E series (Ramesh et al., 2022; Betker et al., 2023), Imagen (Saharia et al., 2022), and Stable Diffusion (Podell et al., 2023), have significantly impacted various domains such as entertainment, design, and education, by enabling high-quality image generation.
+These technologies not only advance the field of text-to-image generation but also bloom applications such as video generation (Blattmann et al., 2023; Zhang et al., 2023d; Tan et al., 2024b), image editing (Song et al., 2021; Huang et al., 2023b; Zhang et al., 2023c) and human image generation (Wang et al., 2024).
+Despite achieving incredible progress, the evaluation methods in this area are far from flawless and suffer heavily from data bias, as they are mainly trained on real images but are employed to evaluate synthesized images.
+Since human-based evaluations are considerably costly in money and time, existing evaluation methods are primarily based on pretrained models, which are trained on real images.
+However, the trained real images are generated by humans and high in image faithfulness and text-image alignment because of their generation essence.
+Meanwhile, the evaluated images are synthesized by text-to-image models and encounter problems such as low image faithfulness or text-image alignment, constrained by the performance of generative models.
+We dub the gap between the training data and the evaluated data as data bias, which may cause the evaluation models perform ill-suited on text-to-image evaluation.
+Because of the data bias, existing text-to-image evaluation methods performs poorly in synthesized image evaluations.
+Unfortunately, during our preliminary observation, nearly every synthesized images contain visual elements with low image faithfulness or text-image alignment, emphasize their significance on evaluation performance.
+Notably, there are also some works such as HPSv2 (Wu et al., 2023b) and PickScore Kirstain et al.
+(2024), where their evaluation models are trained synthesized images.
+However, in their evaluation settings, the utilized synthesized images are treated as real images as they don’t explicitly recognize the problem of synthesized images with low image faithfulness.
+In view of these issues, we propose EvalAlign, a comprehensive, fine-grained and interpretable metric on text-to-image model assessing with low cost but high accuracy.
+To build EvalAlign, we first curate a dataset composed of fine-grained human feedback scores on synthesized images, with consideration of the corresponding prompts.
+The granularity of the feedback covers 11 skills categorized into two aspects: image faithfulness and text-image alignment.
+After that, we Supervised finetune (SFT) a Multimodal Large Language Model (MLLM) on the annotated dataset, aligning it with human prior on detailed and accurate text-to-image evaluation.
+Owing to extensive pre-training and large model capacity, MLLMs demonstrate excellent image-text understanding and generalization capabilities.
+However, since the pre-training data does not include synthesized images with low image faithfulness or evaluation-related text instructions, using MLLMs directly for model evaluation may yield non-optimal results.
+Especially, we want to use MLLMs to support comprehensive and detailed evaluations, encompassing 11 skills and 2 aspects.
+The definitions and nuances of these may not be fully understood by the MLLM.
+Therefore, we employ SFT on a small amount of high-quality annotated data to align the MLLM with human judgement on evaluating synthesized images in criteria of 11 skills and 2 aspects.
+Notably, since the main intelligence of EvalAlign stems from the annotated dataset and the utilized MLLM, we will make them accessible to the public.
 In summary, our main contributions can be summarized as follows:
-We introduce a dataset specifically designed to address the evaluation challenges of text-to-image models.
-This dataset, derived from multiple data sources, has been thoroughly cleaned and systematically annotated by human.
-It enables precise evaluation of text-image alignment and image faithfulness.
-We propose the EvalAlign evaluation metric, which accurately aligns evaluation metrics with human preferences.
-This method is cost-effective in terms of annotation and training, computationally efficient, and provides interpretable metrics.
-We conduct evaluations over 24 text-to-image models and compare EvalAlign with existing evaluation methods.
-Extensive experiments demonstrate that EvalAlign outperforms other methods in evaluating model performance.
+We build a detailed human feedback dataset specifically designed to address the aforementioned challenges of text-to-image model evaluations.
+The annotated dataset is thoroughly cleaned, carefully balanced in topics, and systematically annotated by human.
+The dataset is composed by fine-grained human prior on evaluating synthesized images in criteria of 11 skills and 2 aspects.
+We propose EvalAlign, a text-to-image evaluation method which accurately aligns evaluation models with fine-grained human prior using the annotated dataset.
+EvalAlign exclusively supports an accurate, comprehensive, fine-grained and interpretable text-to-image evaluations.
+Besides EvalAlign is cost-effective in terms of annotation and training and computationally efficient.
+With EvalAlign, we conduct evaluations over 24 text-to-image models and compare EvalAlign with existing evaluation methods.
+Quantitative and qualitative experiments demonstrate that EvalAlign outperforms other methods in evaluating model performance.
 2 Related Work
-Evaluations and benchmarks of text-to-image generation.
-Despite the incredible progress achieved by text-to-image generation, evaluations and benchmarks in this area are far from flawless and contain critical limitations.
-For example, the most commonly used metrics, IS [44], FID [15], and CLIPScore [14] are broadly recognized as inaccurate for their inconsistency with human perception.
-To address, HPS series [55, 54], PickScore [20], and ImageReward [56] introduced human preference prior on image assessing to the benchmark, thereby allowing better correlation with image quality.
+2.1 Benchmarks of text-to-image generation
+Despite the incredible progress achieved by text-to-image generation Zhang et al.
+(2023a); Tan et al.
+(2024a), evaluations and benchmarks in this area are far from flawless and contain critical limitations.
+For example, the most commonly used metrics, IS (Salimans et al., 2016), FID (Heusel et al., 2017), and CLIPScore (Hessel et al., 2021) are broadly recognized as inaccurate for their inconsistency with human perception.
+To address, HPS series (Wu et al., 2023b; a), PickScore (Kirstain et al., 2024), and ImageReward (Xu et al., 2024) introduced human preference prior on image assessing to the benchmark, thereby allowing better correlation with image quality.
 However, with varying source and size of training data, these methods merely score the evaluated images in a coarse and general way, which cannot serve as an indication for model evolution.
-Meanwhile, HEIM [21] combined automatic and human evaluation and holistically evaluated text-to-image generation in 12 aspects, such as alignment, toxicity, and so on.
+Meanwhile, HEIM (Lee et al., 2024) combined automatic and human evaluation and holistically evaluated text-to-image generation in 12 aspects, such as alignment, toxicity, and so on.
 As a consequence, HEIM relies heavily on human labour, limiting its application within budget-limited research groups severely.
-[30] standardized the protocol and settings of human evaluation, ensuring its verifiable and reproducible.
+Otani et al.
+(2023) standardized the protocol and settings of human evaluation, ensuring its verifiable and reproducible.
+Considering the issues of existing benchmarks, we propose EvalAlign to offer a cost-efficient, comprehensive and fine-grained text-to-image model evaluation.
+Through our observations, we found that image faithfulness and text-image alignment are two key factors for comprehensive evaluation.
+Image faithfulness requires the model to generate visual elements that are consistently faithful to the real-world.
+For example, visual elements such as distorted body.
+Meanwhile, text-image alignment measures how the generated images are aligned with their corresponding prompts.
 There are also some works bear a resemble with us.
-For instance, TIFA [16], Gecko [53] and LLMScore [27] also formulate the evaluation as a set of visual question answering procedure and use LLMs as evaluation models.
+For instance, TIFA (Hu et al., 2023), Gecko (Wiles et al., 2024) and LLMScore (Lu et al., 2024) also formulate the evaluation as a set of visual question answering procedure and use LLMs as evaluation models.
 However, while they all mainly focus on text-image alignment, our approach takes both text-image alignment and image faithfulness into consideration.
@@
 Human-aligned | Fine-grained | Automatic | P | I | A | Faithfulness | Alignment | ||
-Inception Score [44] | NeurIPS 2016 | ✗ | ✗ | ✓ | – | 1.3M | – | ✓ | ✗ |
-FID [15] | NeurIPS 2017 | ✗ | ✗ | ✓ | – | 1.3M | – | ✓ | ✗ |
-CLIP-score [14] | EMNLP 2021 | ✗ | ✗ | ✓ | 400M | 400M | – | ✗ | ✓ |
-HPS [55] | ICCV 2023 | ✓ | ✗ | ✓ | 25K | 98K | 25K | – | – |
-TIFA [16] | ICCV 2023 | ✓ | ✓ | ✓ | 4K | – | 25K | ✗ | ✓ |
-TVRHE [30] | CVPR 2023 | ✓ | ✗ | ✗ | – | – | – | ✓ | ✗ |
-ImageReward [56] | NeurIPS 2023 | ✓ | ✗ | ✓ | 8.8K | 68K | 137K | – | – |
-PickScore [20] | NeurIPS 2023 | ✓ | ✗ | ✓ | 35K | 1M | 500K | – | – |
-HPS v2 [54] | arXiv 2023 | ✓ | ✗ | ✓ | 107K | 430K | 645K | – | – |
-HEIM [21] | NeurIPS 2023 | ✓ | ✓ | ✗ | – | – | – | ✓ | ✓ |
-Gecko [53] | arXiv 2024 | ✓ | ✓ | ✓ | 2K | – | 108K | ✗ | ✓ |
-LLMScore [27] | arXiv 2024 | ✓ | ✓ | ✓ | – | – | – | ✗ | ✓ |
+Inception Score (Salimans et al., 2016) | NeurIPS 2016 | ✗ | ✗ | ✓ | – | 1.3M | – | ✓ | ✗ |
+FID (Heusel et al., 2017) | NeurIPS 2017 | ✗ | ✗ | ✓ | – | 1.3M | – | ✓ | ✗ |
+CLIP-score (Hessel et al., 2021) | EMNLP 2021 | ✗ | ✗ | ✓ | 400M | 400M | – | ✗ | ✓ |
+HPS (Wu et al., 2023b) | ICCV 2023 | ✓ | ✗ | ✓ | 25K | 98K | 25K | – | – |
+TIFA (Hu et al., 2023) | ICCV 2023 | ✓ | ✓ | ✓ | 4K | – | 25K | ✗ | ✓ |
+TVRHE (Otani et al., 2023) | CVPR 2023 | ✓ | ✗ | ✗ | – | – | – | ✓ | ✗ |
+ImageReward (Xu et al., 2024) | NeurIPS 2023 | ✓ | ✗ | ✓ | 8.8K | 68K | 137K | – | – |
+PickScore (Kirstain et al., 2024) | NeurIPS 2023 | ✓ | ✗ | ✓ | 35K | 1M | 500K | – | – |
+HPS v2 (Wu et al., 2023a) | arXiv 2023 | ✓ | ✗ | ✓ | 107K | 430K | 645K | – | – |
+HEIM (Lee et al., 2024) | NeurIPS 2023 | ✓ | ✓ | ✗ | – | – | – | ✓ | ✓ |
+Gecko (Wiles et al., 2024) | arXiv 2024 | ✓ | ✓ | ✓ | 2K | – | 108K | ✗ | ✓ |
+LLMScore (Lu et al., 2024) | arXiv 2024 | ✓ | ✓ | ✓ | – | – | – | ✗ | ✓ |
 EvalAlign (ours) | – | ✓ | ✓ | ✓ | 3K | 21K | 132K | ✓ | ✓ |
-Multimodal Large Language Models (MLLMs).
-Pretrained on massive text-only and image-text data, MLLMs have exhibited exceptional image-text joint understanding and generalization abilities, facilitating a large spectrum of downstream applications.
-Among the works major in MLLMs, LLaVA [26, 24] and MiniGPT4 [61, 5] proposed to conduct visual instruction tuning during SFT, so that MLLMs can be easily aligned to human preference and precisely answer fine-grained questions on visual content.
-Meanwhile, Video-LLaMA [58] and VideoChat [22] utilized MLLMs for video understanding.
-VILA [23] quantitatively proved that involving text-only instruction-tuning data during SFT can further ameliorate model performance on text-only and multimodal downstream tasks.
-LLaVA-NeXT [25] extracted visual tokens for both the resized input image and the segmented sub-images to provide more detailed visual information for MLLMs, achieving significant performance bonus on tasks with high-resolution input images.
-In this paper, we manually annotate a minimal amount of high-quality visual instruction-tuning data to adapt MLLMs to perform human preference aligned and fine-grained evaluation.
-Owing to the generalization and multimodal understanding abilities of MLLMs, our experiments demonstrated that the finetuned MLLMs can accurately assess text-to-image generation models by following instructions regarding various fine-grained criteria of synthesised image.
+2.2 Multimodal Large Language Models (MLLMs)
+Pre-trained on massive text-only and image-text data, MLLMs have exhibited exceptional image-text joint understanding and generalization abilities, facilitating a large spectrum of downstream applications.
+Among the works major in MLLMs, LLaVA (Liu et al., 2024b; 2023) and MiniGPT4 (Zhu et al., 2023; Chen et al., 2023a) observed that multimodal SFT is sufficient to align MLLMs with human preferences and enable them to accurately answer fine-grained questions about visual content.
+Besides, Video-LLaMA (Zhang et al., 2023b) and VideoChat (Li et al., 2023) utilized MLLMs for video understanding.
+VILA (Lin et al., 2023) quantitatively proved that involving text-only instruction-tuning data during SFT can further ameliorate model performance on text-only and multimodal downstream tasks.
+LLaVA-NeXT (Liu et al., 2024a) extracted visual tokens for both the resized input image and the segmented sub-images to provide more detailed visual information for MLLMs, achieving significant performance bonus on tasks with high-resolution input images.
+However, due to the data bias, existing MLLMs cannot perfectly quantify for text-to-image evaluations.
+Thus, we meticulously curate a SFT dataset to align MLLMs with detailed human feedback on synthesized images.
 3 EvalAlign Dataset Construction
-To train, validate and test the effectiveness of our evaluation models, we build EvalAlign benchmark.
-Specifically, EvalAlign dataset is a meticulously annotated collection featuring fine-grained annotations for images generated from text prompts.
+To train, validate and test the effectiveness of our evaluation models, we build EvalAlign dataset.
+Specifically, EvalAlign dataset is a meticulously annotated collection featuring fine-grained annotations for images generated on text conditions.
 This dataset comprises 21k images, each accompanied by detailed instructions.
@@
 Prompt collection.
-To assess the capabilities of our model in terms of text-image alignment and image faithfulness, we collect, filter, and clean prompts from existing evaluation datasets and generated prompts based on LLM.
+To assess the capabilities of our model in terms of image faithfulness and text-image alignment, we collect, filter, and clean prompts from existing evaluation datasets and generated prompts based on LLM.
 These prompts encompass a diverse range from real-world user prompts, prompts generated through rule-based templates with LLM, to manually crafted prompts.
-Specifically, the utilized prompts are soureced from HPS [55], HRS-Bench [2], HPSv2 [54], TIFA [16], DSG [7], T2I-Comp [17], Winoground [50], DALL-EVAL [8], DiffusionDB [52], PartiPrompts [57], DrawBench [43], and JourneryDB [49].
+Specifically, the utilized prompts are soureced from HPS (Wu et al., 2023b), HRS-Bench (Bakr et al., 2023), HPSv2 (Wu et al., 2023a), TIFA (Hu et al., 2023), DSG (Cho et al., 2023a), T2I-Comp (Huang et al., 2023a), Winoground (Thrush et al., 2022), DALL-EVAL (Cho et al., 2023b), DiffusionDB (Wang et al., 2023), PartiPrompts (Yu et al., 2022), DrawBench (Saharia et al., 2022), and JourneryDB (Sun et al., 2024).
 Prompt curation.
-The prompts are collected for facilitating fine-grained and comprehensive evaluation on text-to-image models in terms of image faithfulness and text-image alignment.
-Considering some of the collected prompt are unsuitable for these two evaluation tasks, we filter the collected prompts to ensure their quantity, quality and diversity.
-For the image faithfulness evaluation task, we prioritize prompts related to human, animals, and other tangible objects, as prompts depicting sci-fi scenarios are less suitable for this type of assessment.
-Our filtering process initially selects prompts that describe human, animals, and other real objects.
-After deduplicating these prompts, we carefully selected 1,500 distinct prompts with varying topic, background and style for further evaluation.
+To facilitate a clean and reasonable evaluation, each prompt to be annotated have to instruct text-to-image models to generate images that can reflect model performances on image faithfulness and text-image alignment.
+However, considering some of the collected prompts fail to achieve the purpose, we need to filter and balance the collected prompts to ensure their quantity, quality and diversity.
+For image faithfulness evaluation, we prioritize prompts related to human, animals, and other tangible objects, as prompts depicting sci-fi scenarios are less suitable for this type of assessment.
+Consequently, the prompt filter for image faithfulness initially selects prompts that describe human, animals, and other real objects.
+After deduplicating these prompts, we carefully select 1,500 distinct prompts with varying topic, background and style.
 The selected prompts encompass 10k subjects across 15 categories.
-For the text-image alignment task, we refined our selection based on descriptions of style, color, quantity, and spatial relationships in the prompts.
-Only those prompts containing relevant descriptions and exceeding 15 words in length were considered, culminating in a final set of 1,500 prompts for evaluation.
+For text-image alignment evaluation, we refine our selection based on descriptions of style, color, quantity, and spatial relationships in the prompts.
+Specifically, only prompts contain relevant descriptions and exceed 15 words in length are considered, culminating in a final set of 1,500 prompts.
 Image generation.
-To train and evaluate the MLLM, we use a diverse set of images generated by various models using the aforementioned prompts, allowing for detailed human annotation.
+To train and evaluate the MLLM, we use a diverse set of images generated by various models using the aforementioned prompts, facilitating detailed human annotation.
 For each prompt, multiple images are generated across different models.
-The dataset is enriched by images synthesized by models with varying architectures and scales, enhancing the diversity essential for a comprehensive evaluation.
-This variety not only tests the MLLM generalization capabilities but also aids in developing a model with broader applicability.
-The training dataset incorporates images from 8 models, whereas the test dataset spans 24 models.
-The inclusion of 16 unseen models in the test set is crucial for evaluating the MLLM to generalize beyond its training data.
-For detailed information on the inference setting of each model, please refer to the supplementary material.
-This structured approach ensures a robust framework for training and validating the MLLM, positioning it as a versatile and adaptive tool in the field of image generation from textual descriptions.
+The models used to generate these images vary in architectures and scales, enhancing the dataset diversity.
+There are 24 models used to generate these images, varying in architecture as well as scale and thus enhancing the dataset diversity.
+For detailed information on the generation setting of each model, please refer to the appendix.
+The training and validation set comprises synthesized images from 8 out of the 24 models, whereas the test set spans all of them.
+Particularly, the exclusive inclusion of the 16 models in the test set is crucial for validating the MLLM’s ability to generalize beyond its training data.
+Through our manual inspection, in this way, we attain ample synthesized images with a balanced diversity in the performance of image faithfulness and text-image alignment.
 3.2 Data Annotation
@@
 Our annotators extract the entities mentioned in the prompts and label each entity with corresponding attributes, including quantity, color, spatial relationships, and actions.
-During the annotation, we also ask the annotators annotate the overall style of the image if described in the corresponding prompt and report prompts that contain toxic and NSFW content.
+During the annotation, we also ask the annotators to annotate the overall style of the image if described in the corresponding prompt and report prompts that contain toxic and NSFW content.
 These high-quality and detailed annotations facilitate the subsequent SFT training and evaluation of the MLLM.
-This meticulous annotation procedure ensures that the MLLM can accurately align and respond to the nuanced details specified in the prompts, enhancing both the training process and the model’s performance in generating images that faithfully reflect the described attributes and style.
+The prompt annotation procedure ensures that the MLLM can accurately align and respond to the nuanced details specified in the prompts, enhancing both the training process and the model’s performance in generating images that faithfully reflect the described attributes and style.
 Image annotation.
 The images generated by text-to-image models often present challenges such as occluded human body parts, which can impede the effectiveness of SFT training and evaluation of the MLLM.
-Therefore, to address these challenges and enhance the model’s training and evaluative capabilities, specific annotations are applied to all images depicting human and animals.
+To address these challenges and enhance the model’s training and evaluative capabilities, specific annotations are applied to all images depicting human and animals.
 These annotations include: presence of human or animal faces; visibility of hands; visibility of limbs.
@@
 This structured approach to annotation not only aids in identifying common generation errors but also optimizes the model’s ability to generate more accurate and realistic images, thereby improving both training outcomes and the model’s overall performance in generating coherent and contextually appropriate visual content.
-Instruction-Fine-tuning data annotation.
-To align the MLLM with human preference prior on fine-grained image assessing, we can train the model on a minimal amount of fine-grained human preference data through SFT training.
-As a consequence, we devise two sets of questions, each is concentrated on a specific fine-grained aspect of image faithfulness and image-text alignment, and ask human annotators to answer these questions to acquire the fine-grained human preference data.
-To aid the human annotators to understand the meaning of each question and its available answer options, thereby ensuring high annotation quality, we employ a thorough and comprehensive procedure of annotation preparation.
-First, we write a detailed annotation guideline and conduct a training for the annotators to explain the annotation guideline and answer their questions about the annotation (see supplementary materials for the annotation guideline).
+Instruction-finetuning data annotation.
+To align the MLLM with human preference prior on detailed synthesized image assessing, we can train the model on a minimal amount of fine-grained human feedback data through SFT training.
+As a consequence, we devise two sets of questions, each is concentrated on a specific fine-grained skill of image faithfulness and image-text alignment.
+Human annotators are required to answer these questions to acquire the fine-grained human preference data.
+To aid them to understand the meaning and principle of each question, thereby ensuring high annotation quality, we employ a thorough and comprehensive procedure of annotation preparation.
+First, we write a detailed annotation guideline and conduct a training for the annotators to explain the annotation guideline and answer their questions about the annotation.
 Then, we conduct a multi-turn trial annotation on another 50 synthesized images.
@@
 In total, we conduct nine turns of trial annotation, and in the last turn of the trial, the Cohen’s kappa coefficient of our annotators reaches , indicating high inter-annotator reliability and high annotation quality.
-More results about the trial annotation will be included in the supplementary materials.
 After completing the aforementioned preparations, we delegate the images filtered during image annotation to 10 annotators and ask them to complete the annotation just as how they did in the trial annotation.
-Furthermore, during the whole annotation procedure, four text-to-image generation experts conduct random sampling quality inspection on the present annotated results, causing a second and a third re-annotation on 423 and 112 inspection-failed samples.
+Furthermore, during the whole annotation procedure, four experts in text-to-image generation conduct random sampling quality inspection on the present annotated results, causing a second and a third re-annotation on 423 and 112 inspection-failed samples.
 Overall, owing to the valuable work of our human annotators and our fastidious annotation procedure, we get quality-sufficient instruction-tuning data required for the SFT training of the MLLM.
-More details of the annotation procedure will be introduced in the supplementary files.
+More details of the annotation procedure will be introduced in supplementary files.
 3.3 Dataset Statistics
-To summarize, we generate 24k images from 3k prompts based on 8 text-to-image models, which includes DeepFloyd IF [1], SD15 [40], LCM [28], SD21 [40], SDXL [33], Wuerstchen [32], Pixart [6], and SDXL-Turbo [48].
-After data filtering, 4.5k images were selected as annotation data for task of text-image alignment.
-Subsequently, these images were carefully annotated to generate 13.5k text-image pairs, where 11.4k were used to the training dataset and 2.1k to the validation dataset.
-For the image faithfulness task, we select 12k images for annotate, producing 36k text-image pairs, with 30k are used to the training dataset and 6.2k to the validation dataset.
+To summarize, we generate 24k images from 3k prompts based on 8 text-to-image models, which includes DeepFloyd IF (Alex Shonenkov & et al., 2023), SD15 (Rombach et al., 2022a), LCM (Luo et al., 2023), SD21 (Rombach et al., 2022a), SDXL (Podell et al., 2023), Wuerstchen (Pernias et al., 2023), Pixart (Chen et al., 2023b), and SDXL-Turbo (Stability AI, a).
+After data filtering, 4.5k images are selected as annotation data for task of text-image alignment.
+Subsequently, these images are carefully annotated to generate 13.5k text-image pairs, where 11.4k are used to the training dataset and 2.1k to the validation dataset.
+For the image faithfulness task, we select 12k images for annotation, yielding 36k text-image pairs, with 30k are used to the training dataset and 6.2k to the validation dataset.
 Additionally, we employed 24 text-to-image models to generate 2.4k images from 100 prompts.
-After annotation, these images were used as testing dataset.
+After annotation, these images are used as testing dataset.
 Figure 3 and Figure 3 show the distribution of objects in different categories within our prompts, demonstrating the diversity and balance of our prompts.
 4 Training and Evaluation Methods
-4.1 Supervised Fine-Tuning the MLLM
+4.1 Supervised Finetuning the MLLM
 As we mentioned above, we use MLLMs as the evaluation models and let it to answer a set of carefully-designed instructions, thereby achieving quantitative measurement of fine-grained text-to-image generation skills.
-Due to training bias, zero-shot MLLMs perform poorly when it comes to fine-grained evaluation on synthesized images, particularly in term of image faithfulness.
-To address, we conduct SFT training on the annotated data to align the MLLM with human preference prior on fine-grained image assessing.
-Formally, the SFT training sample can be denoted as a triplet, i.e., , where denotes the question (or the instruction), the multimodal input and the answer, respectively.
-Specifically, is a question about a fine-grained aspect of the generated image as well as the available answer options, is mainly the image and necessary textual description, while is the specific answer chosen from the given options and annotated by human.
-Notably, we expand each given option to a more detailed and descriptive version, benefiting the SFT performance by allowing the MLLM to understand the meaning of each option.
-During SFT training, the optimization objective is the original autoregressive loss function utilized to train LLMs, but calculated only on .
-Specifically, the loss function can be formulated as follows, where is the length of the ground truth answer:
+Due to data bias, zero-shot MLLMs perform poorly when it comes to evaluation on generated images, particularly in term of image faithfulness.
+To solve this problem, we apply SFT training on the detailed human annotation to align the MLLM with human preference prior.
+Formally, the SFT training sample can be denoted as a triplet: question (or the instruction), multimodal input and answer.
+During SFT training, the optimization objective is the autoregressive loss function utilized to train LLMs, but calculated only on the answer, the loss function can be formulated as follows:
+where is the length of the ground truth answer, is a fine-grained question of the generated image and its available answer, is the image and textual description, while is the human annotated answer selected from the given options.
+Notably, we expand each option to make it more detailed and descriptive, thereby benefiting SFT performance by allowing the MLLM to better understand the meaning of each option.
 4.2 Evaluation and Metrics
-During inference, the MLLM, parameterized by , is required to generate the appropriate response , given the specific question and the multimodal input in an autoregressive way:
-This autoregressive generation procedure is considered complete once the model generates a EOS token or the generated response exceeds the preset maximum generation length.
-After the generation, we use rule-based filtering and regular expression to extract the option chosen by the MLLM, where each option has an exclusive predefined score to quantitatively measure a fine-grained skill specified by the question :
-where represents the procedure of option extraction and score searching.
+To evaluate synthesized images with consideration of its synthetic nature, EvalAlign is designed to evaluate image faithfulness and text-image alignment in a fine-grained way.
+Notably, image faithfulness and text-image alignment are two common errors occurred in synthesized images, whereas real images inherently exhibit high levels of both image faithfulness and text-image alignment.
+Image Faithfulness measures whether synthesized images are faithful to real-world commonsense.
+With higher image faithfulness, the visual elements of generated images more closely resemble their real-world counterparts.
+Unfortunately, text-to-image models often generate images with low faithfulness, such as distorted body structures and human hands.
+This is also a critical reason why we set image faithfulness as one of the benchmarking aspects in EvalAlign.
+Additionally, evaluating image faithfulness requires considering the input prompts, as prompts may describe unreal or impossible scenarios that inherently affect the faithfulness of the generated images.
+For example, when prompts like "a dog walking like a human" or "a man on Mars without a spacesuit" are provided, the generated images may naturally deviate from real-world image faithfulness.
+Under such circumstances, the synthesized images cannot be regarded as low in image faithfulness since the generative models are merely following prompts that contain super-reality scenarios.
+Text-Image Alignment evaluates whether generated images are aligned with their conditioned prompts.
+In the inference settings of text-to-image models, the image generation process is conditioned on textual prompts, requiring alignment between the text prompts and the synthesized images.
+However, through our observations, text-to-image models cannot consistently follow input prompts, often yielding images with visual elements misaligned with the input prompts.
+For example, models may generate images featuring an orange cat when conditioned on the text prompt "a blue cat."
+During inference, the multimodal large language model (MLLM) is required to generate an appropriate response given a specific question and multimodal input in an autoregressive manner:
+where is the -th generated token, represents the sequence of tokens generated before step , and denotes the parameters of the fine-tuned MLLM.
+This autoregressive generation process is considered complete once the model generates an end-of-sequence (EOS) token or the generated response exceeds a preset maximum generation length.
+After generation, we employ rule-based filtering and regular expressions to extract the option chosen by the MLLM.
+Each option is assigned a unique predefined score to quantitatively measure a fine-grained skill specified by the question :
+where represents the procedure of option extraction and score mapping.
 We devise two holistic and detailed question sets, and , that encompass every aspect of image faithfulness and text-image alignment, respectively.
-Consequently, Our metric EvalAlign can be defined by simply summing the scores of the two question sets:
-where and indicates the image faithfulness score and image-text alignment score evaluated by EvalAlign.
+Consequently, our metric, EvalAlign, can be defined by averaging the scores of the questions in the two sets:
+where and indicate the image faithfulness score and the text-image alignment score evaluated by our method, respectively.
 4.3 Implementation Details
-We leverage the LLaVA-NeXT model and applied the LoRA fine-tuning method to align it with the training dataset as described previously.
-The entire training process was conducted on NVIDIA A100 GPUs with hours, using a learning rate of 510-5.
-Additionally, we restricted the LoRA fine-tuning to the Q and K weights of the attention module, as extending the fine-tuning to the ViT and projection modules was found to cause overfitting.
-Importantly, our method does not necessitate any modifications to the MLLM, rendering it compatible with any MLLM model.
-For the ablation study, we evaluated the LLaVA-NeXT 13B model on the validation dataset.
-In the final experiment, we apply supervised fine-tuning to the LLaVA-NeXT 34B model on the testing dataset and compare against different evaluation metrics.
+For details about the SFT training, we apply LoRA (Hu et al., 2021) finetuning on LLaVA-NeXT (Liu et al., 2024a) models to align them with the EvalAlign dataset.
+Additionally, we merely adapt LoRA finetuning on the Q and K weights of the attention module, as extending the finetuning to the ViT (Dosovitskiy, 2020) and projection modules will lead to overfitting.
+The entire training process is conducted on NVIDIA A100 GPUs for hours, with a learning rate of 510-5.
+As for the ablation study, we evaluate the finetuned LLaVA-NeXT 13B model on the validation dataset.
+In the final experiment, we apply SFT to the LLaVA-NeXT 34B model on the testing dataset to testify its generalization ability.
 5 Experimental Results
-In this section, we show results of EvalAlign on benchmarking 24 text-to-image generative models.
-We also compare with popular text-to-image metrics and frameworks, including Inception Score (IS) [44], Frechet Inception Distance (FID) [15], CLIP-score [14], and HPS v2 [54].
-Please refer to supplementary materials for detailed results and analyses.
 5.1 Main Results
-Model | Human | EvalAlign |
-PixArt XL2 1024 MS [6] | 2.2848 | 1.6415 |
-Dreamlike Photoreal v2.0 [11] | 2.0070 | 1.4522 |
-SDXL Refiner v1.0 [47] | 1.9229 | 1.6072 |
-SDXL v1.0 [33] | 1.8136 | 1.4675 |
-Wuerstchen [32] | 1.7837 | 1.4279 |
-LCM SDXL [28] | 1.6910 | 1.3391 |
-Openjourney [34] | 1.6667 | 1.175010 |
-Safe SD MAX [31] | 1.6491 | 1.2175 |
-LCM LORA SDXL [28] | 1.6387 | 1.3833 |
-Safe SD STRONG [31] | 1.630810 | 1.146611 |
-Safe SD MEDIUM [31] | 1.627511 | 1.129815 |
-Safe SD WEAK [31] | 1.607812 | 1.118817 |
-SD v2.1 [40] | 1.552413 | 1.109418 |
-SD v2.0 [40] | 1.527714 | 1.130014 |
-Openjourney v2 [35] | 1.500015 | 0.995620 |
-Redshift diffusion [39] | 1.473316 | 1.138212 |
-Dreamlike Diffusion v1.0 [10] | 1.465217 | 1.2052 |
-SD v1.5 [40] | 1.441718 | 1.136213 |
-IF-I-XL v1.0 [1] | 1.380819 | 0.922122 |
-SD v1.4 [40] | 1.359220 | 0.951121 |
-Vintedois Diffusion v0.1 [51] | 1.356221 | 1.079719 |
-IF-I-L v1.0 [1] | 1.263522 | 0.881423 |
-MultiFusion [29] | 1.237223 | 1.129816 |
-IF-I-M v1.0 [1] | 1.013524 | 0.792824 |
-Model | Human | EvalAlign |
-IF-I-XL v1.0 [1] | 5.4500 | 5.5300 |
-IF-I-L v1.0 [1] | 5.2300 | 5.4500 |
-SDXL Refiner v1.0 [47] | 5.2100 | 5.4000 |
-LCM SDXL [28] | 5.1800 | 5.3300 |
-PixArt XL2 1024 MS [6] | 5.1100 | 5.3100 |
-IF-I-M v1.0 [1] | 5.0800 | 5.2200 |
-LCM LORA SDXL [28] | 5.0600 | 5.2700 |
-SDXL v1.0 [33] | 5.0300 | 5.3500 |
-Wuerstchen [32] | 4.8700 | 5.1700 |
-Openjourney [34] | 4.830010 | 4.920015 |
-SD v2.1 [40] | 4.800011 | 5.070011 |
-MultiFusion [29] | 4.680012 | 4.800018 |
-Dreamlike Diffusion v1.0 [10] | 4.660013 | 5.150010 |
-SD v2.0 [40] | 4.640014 | 5.010012 |
-Vintedois Diffusion v0.1 [51] | 4.620015 | 4.950014 |
-Safe SD STRONG [31] | 4.600016 | 4.830017 |
-Dreamlike Photoreal v2.0 [11] | 4.560017 | 4.980013 |
-Safe SD WEAK [31] | 4.530018 | 4.710020 |
-SD v1.4 [40] | 4.520019 | 4.760019 |
-SD v1.5 [40] | 4.450020 | 4.900016 |
-Safe SD MEDIUM [31] | 4.400021 | 4.560024 |
-Redshift diffusion [39] | 4.350022 | 4.670021 |
-Safe SD MAX [31] | 4.310023 | 4.590023 |
-Openjourney v2 [35] | 4.150024 | 4.650022 |
-Evaluation on image faithfullness.
-We conducted the image faithfullness evaluation on the testing dataset, as detailed in Table 3.
-Our evaluation metric, EvalAlign, demonstrates close alignment with human preferences Specifically, the rankings of the top and bottom 10 models by both EvalAlign and human evaluation scores show remarkable consistency, confirming that our metric closely mirrors human evaluation.
-Importantly, most of the images from the 24 generative models were not present during training, showcasing the robust generalization capability of our evaluation method.
+Model | Human | EvalAlign | HPS v2 | CLIP-score | ImageReward | PickScore |
+PixArt XL2 1024 MS (Chen et al., 2023b) | 2.2848 | 1.6415 | 31.6226 | 0.8580 | 0.9696 | 22.1335 |
+Dreamlike Photoreal v2.0 (dreamlike.art, b) | 2.0070 | 1.4522 | 29.2322 | 0.828612 | 0.1886 13 | 21.2271 |
+SDXL Refiner v1.0 (Stability AI, b) | 1.9229 | 1.6072 | 29.8197 | 0.8566 | 0.7245 | 22.0492 |
+SDXL v1.0 (Podell et al., 2023) | 1.8136 | 1.4675 | 29.0620 | 0.8467 | 0.7043 | 21.8106 |
+Wuerstchen (Pernias et al., 2023) | 1.7837 | 1.4279 | 30.6622 | 0.819914 | 0.3212 11 | 21.3720 |
+LCM SDXL (Luo et al., 2023) | 1.6910 | 1.3391 | 29.3588 | 0.833510 | 0.5304 | 21.6532 |
+Openjourney (PromptHero, a) | 1.6667 | 1.175010 | 26.347513 | 0.819615 | 0.1478 16 | 20.863710 |
+Safe SD MAX (Patrick et al., 2022) | 1.6491 | 1.2175 | 25.739617 | 0.755524 | -0.050722 | 20.459421 |
+LCM LORA SDXL (Luo et al., 2023) | 1.6387 | 1.3833 | 27.329910 | 0.8364 | 0.4959 | 21.4824 |
+Safe SD STRONG (Patrick et al., 2022) | 1.630810 | 1.146611 | 25.576418 | 0.816518 | -0.102223 | 20.621118 |
+Safe SD MEDIUM (Patrick et al., 2022) | 1.627511 | 1.129815 | 26.279814 | 0.810120 | 0.2042 12 | 20.788012 |
+Safe SD WEAK (Patrick et al., 2022) | 1.607812 | 1.118817 | 26.118015 | 0.780923 | -0.126424 | 20.387324 |
+SD v2.1 (Rombach et al., 2022a) | 1.552413 | 1.109418 | 26.582312 | 0.8377 | 0.4116 | 21.0502 |
+SD v2.0 (Rombach et al., 2022a) | 1.527714 | 1.130014 | 25.348121 | 0.817017 | 0.0872 18 | 20.752913 |
+Openjourney v2 (PromptHero, b) | 1.500015 | 0.995620 | 24.698423 | 0.795822 | -0.041521 | 20.408822 |
+Redshift diffusion (Redshift-Diffusion, ) | 1.473316 | 1.138212 | 25.157222 | 0.810121 | 0.0218 20 | 20.615519 |
+Dreamlike Diffusion v1.0 (dreamlike.art, a) | 1.465217 | 1.2052 | 29.6506 | 0.8543 | 0.6508 | 21.2664 |
+SD v1.5 (Rombach et al., 2022a) | 1.441718 | 1.136213 | 25.497219 | 0.821413 | 0.1686 14 | 20.714316 |
+IF-I-XL v1.0 (Alex Shonenkov & et al., 2023) | 1.380819 | 0.922122 | 27.4512 | 0.8449 | 0.6087 | 20.747414 |
+SD v1.4 (Rombach et al., 2022a) | 1.359220 | 0.951121 | 25.369720 | 0.819016 | 0.1050 17 | 20.653517 |
+Vintedois Diffusion v0.1 (Vintedois-Diffusion v0., 1) | 1.356221 | 1.079719 | 26.590111 | 0.8341 | 0.3562 10 | 20.835811 |
+IF-I-L v1.0 (Alex Shonenkov & et al., 2023) | 1.263522 | 0.881423 | 27.4836 | 0.8384 | 0.4463 | 20.717015 |
+MultiFusion (Marco et al., 2023) | 1.237223 | 1.129816 | 23.813324 | 0.815119 | 0.0695 19 | 20.478020 |
+IF-I-M v1.0 (Alex Shonenkov & et al., 2023) | 1.013524 | 0.792824 | 25.952216 | 0.832911 | 0.1637 15 | 20.403523 |
+Model | Human | EvalAlign | HPS v2 | CLIP-score | ImageReward | PickScore |
+IF-I-XL v1.0 (Alex Shonenkov & et al., 2023) | 5.4500 | 5.5300 | 32.547710 | 0.8579 | 0.4391 | 21.199810 |
+IF-I-L v1.0 (Alex Shonenkov & et al., 2023) | 5.2300 | 5.4500 | 32.7140 | 0.8538 | 0.3820 | 21.128412 |
+SDXL Refiner v1.0 (Stability AI, b) | 5.2100 | 5.4000 | 35.6465 | 0.8528 | 0.4738 | 22.3532 |
+LCM SDXL (Luo et al., 2023) | 5.1800 | 5.3300 | 33.8011 | 0.8512 | 0.3833 | 21.9620 |
+PixArt XL2 1024 MS (Chen et al., 2023b) | 5.1100 | 5.3100 | 37.0493 | 0.8634 | 0.6542 | 22.3926 |
+IF-I-M v1.0 (Alex Shonenkov & et al., 2023) | 5.0800 | 5.2200 | 31.095114 | 0.8434 | 0.0499 10 | 20.827020 |
+LCM LORA SDXL (Luo et al., 2023) | 5.0600 | 5.2700 | 32.7752 | 0.834910 | 0.1618 | 21.7627 |
+SDXL v1.0 (Podell et al., 2023) | 5.0300 | 5.3500 | 35.1593 | 0.8540 | 0.4322 | 22.1291 |
+Wuerstchen (Pernias et al., 2023) | 4.8700 | 5.1700 | 36.4632 | 0.8381 | 0.2513 | 21.7779 |
+Openjourney (PromptHero, a) | 4.830010 | 4.920015 | 31.149512 | 0.817316 | -0.086714 | 21.116313 |
+SD v2.1 (Rombach et al., 2022a) | 4.800011 | 5.070011 | 31.101713 | 0.827814 | -0.045312 | 21.2093 |
+MultiFusion (Marco et al., 2023) | 4.680012 | 4.800018 | 28.795724 | 0.826415 | -0.133715 | 20.962517 |
+Dreamlike Diffusion v1.0 (dreamlike.art, a) | 4.660013 | 5.150010 | 34.8196 | 0.8493 | 0.2295 | 21.5550 |
+SD v2.0 (Rombach et al., 2022a) | 4.640014 | 5.010012 | 30.615317 | 0.829813 | -0.142416 | 21.190511 |
+Vintedois Diffusion v0.1 (Vintedois-Diffusion v0., 1) | 4.620015 | 4.950014 | 31.950311 | 0.831912 | -0.022211 | 21.114114 |
+Safe SD STRONG (Patrick et al., 2022) | 4.600016 | 4.830017 | 30.661516 | 0.775123 | -0.502822 | 20.749121 |
+Dreamlike Photoreal v2.0 (dreamlike.art, b) | 4.560017 | 4.980013 | 33.7712 | 0.834411 | -0.085913 | 21.4832 |
+Safe SD WEAK (Patrick et al., 2022) | 4.530018 | 4.710020 | 30.564418 | 0.814018 | -0.272818 | 20.989916 |
+SD v1.4 (Rombach et al., 2022a) | 4.520019 | 4.760019 | 29.914920 | 0.804820 | -0.343819 | 20.846219 |
+SD v1.5 (Rombach et al., 2022a) | 4.450020 | 4.900016 | 30.167319 | 0.814217 | -0.221317 | 20.864018 |
+Safe SD MEDIUM (Patrick et al., 2022) | 4.400021 | 4.560024 | 30.782015 | 0.797421 | -0.359120 | 21.025715 |
+Redshift diffusion (Redshift-Diffusion, ) | 4.350022 | 4.670021 | 29.286522 | 0.806619 | -0.417221 | 20.632723 |
+Safe SD MAX (Patrick et al., 2022) | 4.310023 | 4.590023 | 29.812621 | 0.760124 | -0.609524 | 20.704622 |
+Openjourney v2 (PromptHero, b) | 4.150024 | 4.650022 | 29.238923 | 0.785122 | -0.605123 | 20.597324 |
+Evaluation on image faithfulness.
+We evaluate image faithfulness on the testing dataset to ensure that the finetuned MLLM aligns with human judgment and generalizes to unseen data.
+As detailed in Table 2, the finetuned MLLM successfully aligns with human preferences on image faithfulness, indicating its ability of image faithfulness evaluation is close to human.
+Specifically, the rankings of the top and bottom 10 models by both EvalAlign and human evaluation scores are remarkably consistent.
+Besides, most of the images in the testing dataset, especially those from the 16 exclusive generative models, are not present during the SFT training, showcasing the robust generalization capability of our models.
 Evaluation on text-image alignment.
-The evaluation of text-image alignment on the testing dataset was conducted follow the same manner of image faithfullness evaluation.
-Table 3 reveals that the rankings based on the EvalAlign metric and human evaluation metric for the 24 models are generally consistent.
-This consistency likely stems from the proximity of MLLM’s pre-training tasks to the text-image alignment evaluation tasks, resulting in better performance on text-image alignment evaluation task.
+The evaluation of text-image alignment on the testing dataset is similar to that of image faithfulness.
+Table 2 reveals that the rankings of the 24 evaluated models by EvalAlign are generally consistent with human annotators.
+We believe that the consistency on image faithfulness and text-image alignment evaluations mainly stems from our annotated high-quality SFT dataset.
+It also proves that, with the annotated dataset and the extraordinary image-text joint understanding ability owned by MLLMs, we can easily finetune a MLLM to conduct the evaluation with low cost but close-to-human performance.
 5.2 Ablations and Analyses of EvalAlign
 Results on different prompt categories.
-SFT significantly enhances the performance of MLLMs on evaluation tasks.
-We conducted experiments comparing the LLava-Next 13B model with and without SFT.
-As shown in Table 5 and Table 5, the results demonstrate that SFT considerably improves performance across all prompt categories in both image faithfulness and text-to-image alignment, closely aligning the MLLM’s predictions with human evaluations.
-Note that Table 5 shows that the baseline method without SFT performs poorly on the image faithfulness task, which suggests the necessity of SFT.
+Since MLLMs are not specifically trained to perform evaluations, they are naturally ill-suited for this task, hindering their task performances.
+Therefore, we need to annotate SFT data for this task and finetune the MLLMs accordingly.
+To verify the necessity, We conduct experiments comparing the LLava-Next 13B model with and without SFT.
+As shown in Table 5 and Table 5, the results demonstrate that SFT training considerably improves performance across all prompt categories in both image faithfulness and text-to-image alignment, closely aligning the MLLM’s predictions with human evaluations.
+Note that Table 5 illustrates that the baseline method without SFT performs poorly in image faithfulness and text-image alignment evaluations, particularly in the former.
 Method | Body | Hand | Face | Object | Common |
@@
 34B | 2.1131 | 1.8621 | 1.8083 | 1.3906 | 1.3076 | 1.3921 | 1.2037 | 1.0143 |
-Effect of dataset size for vision-language model training.
-In order to explore the effects of data size, we train model on image faithfulness evaluation task under three different data size: 200, 500 and 800.
-As illustrated in Table 6, performance generally enhances with more training data.
-Notably, training with just 500 data nearly maximizes accuracy, with further increases to 800 data yielding only slight improvements.
-This result suggests that our method requires only a small amount of annotated data to achieve good performance, highlighting the cost-effectiveness of our approach.
+Effect of training dataset size for vision-language model training.
+In order to explore the effects of data size and determine the sufficient amount of training data, we train the model on image faithfulness evaluation task with images and their annotations sourced from 200, 500 and 800 prompts.
+As illustrated in Table 6, the evaluation performance continuously enhances as more training data is used.
+Notably, training with just 500 prompts nearly maximizes accuracy, with further increases to 800 data yielding only marginal improvements.
+This result suggests that our method requires only a small amount of annotated data to achieve good performance, highlighting its cost-effectiveness.
+Generally, since more data leads to better performance, we use all of the available data to finetune our models and release this data to the research community to bootstrap further study.
 Effect of model size.
-Transformers have demonstrated a strong capacity to scale effectively across a variety of tasks, as noted in key studies [9, 36].
-In Table 7, we illustrate the advantages of scaling up the MLLM for image faithfulness evaluation.
-Specifically, increasing the MLLM model size from 7B to 34B results in notable improvements in performance.
+Since transformers are known for their scalability (Radford et al., 2018; Dehghani et al., 2023), we investigate the effect of the model size on the performance of image faithfulness evaluation.
+As illustrated in Table 7, the benefits of scaling up the utilized MLLMs are remarkably significant, where increasing the model size from 7B to 34B results in substantial improvements in evaluation performance.
+For this consequence, for the final version of the EvalAlign evaluation model, we choose LLaVA-NeXT 34B, the largest model in LLaVA-NExT series, and finetune it on our meticulously curated SFT data.
+Since some users of EvalAlign cannot afford MLLM inference with 34B parameters, we will make the 13B and 34B models publicly available.
 5.3 Comparison with Existing Evaluation Methods
@@
 However, the alignment of the vanilla MLLM is relatively low due to the absence of images generated by model (such as distorted bodies and hands images) and issues related to evaluation in the MLLM’s pre-training dataset.
-After applying SFT on the LLaVA-Next 13B model using human annotated data, the model’s predictions on various fine-grained evaluation metrics are almost align to the human-annotated data and significantly surpass the evaluation results of all MLLM models that were not fine-tuned.
+After applying SFT on the LLaVA-Next 13B model using human annotated data, the model’s predictions on various fine-grained evaluation metrics are almost align to the human-annotated data and significantly surpass the evaluation results of all MLLM models that are not finetuned.
 This experimental results confirms that our SFT training enables the MLLM to be successfully applied to the task of evaluating text-to-image models.
-Comparison with state-of-the-art methods.
-To verify the human preference alignment of our model, especially when compared with other baseline methods, we calculate Kendall rank [19] and Pearson [12] correlation coefficient and summarize the results in Table 8.
-This comparison experiment encompasses images generated by 24 text-to-image models, covering an extensive range of synthesized image distribution.
+Comparison with other methods.
+To verify the human preference alignment of our model, especially when compared with other baseline methods, we calculate Kendall rank (KENDALL, 1938) and Pearson (Freedman et al., 2007) correlation coefficient on images generated by 24 text-to-image models and summarize the results in Table 8.
 Method | Faithfulness | Alignment | ||
@@
 6 Conclusion and Discussion
-In this work, we design an economic and efficient evaluation method that offers high accuracy, strong generalization capabilities, and provides fine-grained, interpretable metrics.
-We develop a comprehensive data annotation and cleaning process tailored for evaluation tasks, and established the EvalAlign benchmark for training and evaluating models on supervised fine-tuning tasks for MLLMs.
+In this work, we design an economic evaluation method that offers high accuracy, strong generalization capabilities, and provides fine-grained, interpretable metrics.
+We develop a comprehensive data annotation and cleaning process tailored for evaluation tasks, and establish the EvalAlign benchmark for training and evaluating models on supervised fine-tuning tasks for MLLMs.
 Experimental results across 24 text-to-image models demonstrate that our evaluation metrics surpass the accuracy of all the state-of-art evaluation method.
@@
 We hope that our work can inspire and facilitate future research in this field.
+7 Reproducibility Statement
+The full version of the source code, dataset, as well as the final version of the finetuned MLLMs (one finetuned on LLaVA-NeXT 13B and the other one finetuned on LLaVA-NeXT 34B) will be released to the public.
+The data construction procuedure, including data collection and curation, data cleaning and annotation, is thoroughly described in Section 3.
+For details related to the human annotation and the measures that used to ensure its quality, we comprehensively introduce them in Appendix B.
+As for every experiment introduced in this paper, we provide a general introduction in Section 5 and exhibit implementation details related to reproduce our experiments.
+Specifically, the latter includes the hyper-parameters of each evaluated models, the employed instruction, as well as more supplementary experiments, which are described in Appendix C, Appendix D and Appendix E.
+8 Ethics Statement
+We are committed to conducting this research with the highest ethical standards.
+Our goal is to contribute positively to the fields of evaluation benchmarks on artificial intelligence generated content, emphasizing transparency and reproducibility in our design.
+Similar with other MLLMs, EvalAlign may potentially generate responses contain offensive, inappropriate, or harmful content.
+Since the base MLLMs of EvalAlign are pretrained on large datasets scraped from the web that might contain private information and harmful content, they may inadvertently generate or expose sensitive information, raising ethical and privacy concerns.
+MLLMs are also susceptible to adversarial attacks, where inputs are intentionally crafted to deceive the model.
+This vulnerability can be exploited to manipulate model outputs, posing security and ethic risks.
+To alleviate these safety limitation and our fulfill our social responsibility as artificial intelligence researchers, we create dedicated evaluation sets for bias detection and mitigation, and conducted adversarial testing through hours of redteaming.
+Besides, EvalAlign is designed for fine-grained, human-aligned automatic text-to-image evaluations, which can serve as a stepping stones toward revealing the inner generation nature of text-to-image generative models, thereby lowering the ethical hazard of these models.
+We believe that with appropriate use, it could provide users with interesting experiences for detailed synthesized image evaluation, and inspires more appealing research works about text-to-image generation.
 Appendix A Limitations
@@
 Furthermore, human annotation can be time-consuming and resource-intensive, limiting the scale at which we can afford.
-Appendix B Ethical and Social Impacts
-Similar with other MLLMs, EvalAlign may potentially generate responses contain offensive, inappropriate, or harmful content.
-Since the base MLLMs of EvalAlign are pretrained on large datasets scraped from the web that might contain private information and harmful content, they may inadvertently generate or expose sensitive information, raising ethical and privacy concerns.
-MLLMs are also susceptible to adversarial attacks, where inputs are intentionally crafted to deceive the model.
-This vulnerability can be exploited to manipulate model outputs, posing security and ethic risks.
-To alleviate these safety limitation, we create dedicated evaluation sets for bias detection and mitigation, and conducted adversarial testing through hours of redteaming.
-Besides, EvalAlign is designed for fine-grained, human-aligned automatic text-to-image evaluations.
-We believe that with appropriate use, it could provide users with interesting experiences for detailed synthesized image evaluation, and inspires more appealing research works about text-to-image generation.
-Appendix C Annotation Details
+Appendix B Annotation Details
 Before performing the final human annotation, we made a series of efforts to guarantee its quantity, quality and efficiency.
@@
 As for the last two turns of quality inspection, they both revealed zero failed samples due to the thoughtful and rigorous annotation preparation.
-Appendix D Additional Details of the Evaluated Models
+Appendix C Additional Details of the Evaluated Models
 In this section, we introduce the details of the evaluated text-to-image generative models in this work.
 Stable Diffusion {v1.4, v1.5, v2 base, v2.0, v2.1}.
-Stable Diffusion (SD) is a series of 1B text-to-image generative models based on latent diffusion model [41] and is trained on LAION-5B [45].
+Stable Diffusion (SD) is a series of 1B text-to-image generative models based on latent diffusion model (Rombach et al., 2022b) and is trained on LAION-5B (Schuhmann et al., 2022).
 Specifically, the SD series includes SD v1.1, SD v1.2, SD v1.4, SD v1.5, SD v2 base, SD v2.0, and SD v2.1 respectively.
@@
 It is a Transformer Latent Diffusion Model that uses one fixed, pretrained text encoders (T5)) and one latent feature encoder (VAE).
-Latent Consistency Model Stable Diffusion XL Latent Consistency Model Stable Diffusion XL (LCM SDXL) [28] enables SDXL for swift inference with minimal steps.
+Latent Consistency Model Stable Diffusion XL Latent Consistency Model Stable Diffusion XL (LCM SDXL) Luo et al.
+(2023) enables SDXL for swift inference with minimal steps.
 Viewing the guided reverse diffusion process as solving an augmented probability flow ODE (PF-ODE), LCMs are designed to directly predict the solution of such ODE in latent space, mitigating the need for numerous iterations and allowing rapid, high-fidelity sampling.
-Dreamlike Diffusion 1.0. Dreamlike Diffusion 1.0 [10] is a SD v1.5 model finetuned on high-quality art images by dreamlike.art.
+Dreamlike Diffusion 1.0. Dreamlike Diffusion 1.0 (dreamlike.art, a) is a SD v1.5 model finetuned on high-quality art images by dreamlike.art.
 Dreamlike Photoreal 2.0.
-Dreamlike Photoreal 2.0 [11] is a photorealistic text-to-image latent diffusion model resumed from SD v1.5 by dreamlike art.
+Dreamlike Photoreal 2.0 (dreamlike.art, b) is a photorealistic text-to-image latent diffusion model resumed from SD v1.5 by dreamlike art.
 This model was finetuned on 768x768 images, it works pretty good with resolution 768x768, 640x896, 896x640 and higher resolution such as 768x1024.
-Redshift Diffusion. Redshift Diffusion [39] is a Stable Diffusion model finetuned on high-resolution 3D artworks.
+Openjourney v1, v2.
+Openjourney (PromptHero, a) is an open-source text-to-image generation model resumed from SD v1.5 and finetuned on Midjourney images by PromptHero.
+Openjourney v2 (PromptHero, b) was further finetuned using another 124000 images for 12400 steps, about 4 epochs and 32 training hours.
+Redshift Diffusion. Redshift Diffusion (Redshift-Diffusion, ) is a Stable Diffusion model finetuned on high-resolution 3D artworks.
 Vintedois Diffusion.
-Vintedois Diffusion [51] is a Stable Diffusion v1.5 model finetuned on a large number of high-quality images with simple prompts to generate beautiful images without a lot of prompt engineering.
+Vintedois Diffusion (Vintedois-Diffusion v0., 1) is a Stable Diffusion v1.5 model finetuned on a large number of high-quality images with simple prompts to generate beautiful images without a lot of prompt engineering.
 Safe Stable Diffusion {Weak, Medium, Strong, Max}.
-Safe Stable Diffusion [31] is an enhanced version of the SD v1.5 model by mitigating inappropriate degeneration caused by pretraining on unfiltered web-crawled datasets.
+Safe Stable Diffusion (Patrick et al., 2022) is an enhanced version of the SD v1.5 model by mitigating inappropriate degeneration caused by pretraining on unfiltered web-crawled datasets.
 For instance SD may unexpectedly generate nudity, violence, images depicting self-harm, and otherwise offensive content.
@@
 MultiFusion.
-MultiFusion [29] is a multimodal, multilingual diffusion model that extends the capabilities of SD v1.4 by integrating various modules to transfer capabilities to the downstream model.
+MultiFusion (Marco et al., 2023) is a multimodal, multilingual diffusion model that extends the capabilities of SD v1.4 by integrating various modules to transfer capabilities to the downstream model.
 This combination results in novel decoder embeddings, which enable prompting of the image generation model with interleaved multimodal, multilingual inputs, despite being trained solely on monomodal data in a single language.
 DeepFloyd-IF { M, L, XL } v1.0.
-DeepFloyd-IF [1] is a novel state-of-the-art open-source text-to-image model with a high degree of photorealism and language understanding.
+DeepFloyd-IF (Alex Shonenkov & et al., 2023) is a novel state-of-the-art open-source text-to-image model with a high degree of photorealism and language understanding.
 It is a modular composed of a frozen text encoder and three cascaded pixel diffusion modules: a base model that generates 64x64 image based on text prompt and two super-resolution models, each designed to generate images of increasing resolution: 256x256 and 1024x1024, respectively.
@@
 M has 0.4B parameters, L has 0.9B parameters, and XL has 4.3B parameters.
-Throughout all of the conducted experiments, we use the default inference settings for each evaluated models.
-Appendix E Instruction Templates
+Appendix D Instruction Templates
 Here, we present every instruction used for EvalAlign evaluation on image faithfuleness and text-image alignment.
@@
 For images that do not involve the specified attribute, the corresponding question template is not filled in and subsequently input into EvalAlign.
-Appendix F Additional Quantitative Analysis
-F.1 Generalization Experiments
+Appendix E Additional Quantitative Analysis
+E.1 Generalization Experiments
 To verify the generalization capability of our evaluation model, We compared MLLM’s SFT using different training datasets: one with images generated by all 8 text-to-image models and another with images generated by only 4 of these models, while the final evaluation was conducted on images generated by the other 4 models.
@@
 4 | 1.7832 | 1.3526 | 1.9637 | 1.9876 | 1.6891 | 1.8954 | 0.0469 |
-F.2 Instruction Enhancement Experiments
+E.2 Instruction Enhancement Experiments
 Method | Instruction | SDXL | Pixart | Wuerstchen | SDXL-Turbo | IF | SD v1.5 | SD v2.1 | LCM | MAE |
@@
 Additionally, this approach consistently improved evaluation performance across different text-to-image models.
-F.3 Mulit-scaling Resolutions Experiments
+E.3 Mulit-scaling Resolutions Experiments
 Method | Multi Scale | SDXL | Pixart | Wuerstchen | SDXL-Turbo | IF | SD v1.5 | SD v2.1 | LCM | MAE |
@@
 As shown in Table 13, training with multi-scale inputs significantly enhanced the model’s understanding of image and achieved better evaluation performance.
-F.4 Full Comparison with Existing Evaluation Methods
-Model | Human | EvalAlign | HPS v2 | CLIP-score | ImageReward | PickScore | IS |
-PixArt XL2 1024 MS [6] | 2.2848 | 1.6415 | 31.6226 | 0.8580 | 0.9696 | 22.1335 | 5.258321 |
-Dreamlike Photoreal v2.0 [11] | 2.0070 | 1.4522 | 29.2322 | 0.828612 | 0.1886 13 | 21.2271 | 5.645214 |
-SDXL Refiner v1.0 [47] | 1.9229 | 1.6072 | 29.8197 | 0.8566 | 0.7245 | 22.0492 | 5.398516 |
-SDXL v1.0 [33] | 1.8136 | 1.4675 | 29.0620 | 0.8467 | 0.7043 | 21.8106 | 5.377418 |
-Wuerstchen [32] | 1.7837 | 1.4279 | 30.6622 | 0.819914 | 0.3212 11 | 21.3720 | 5.9751 |
-LCM SDXL [28] | 1.6910 | 1.3391 | 29.3588 | 0.833510 | 0.5304 | 21.6532 | 5.9874 |
-Openjourney [34] | 1.6667 | 1.175010 | 26.347513 | 0.819615 | 0.1478 16 | 20.863710 | 5.736810 |
-Safe SD MAX [31] | 1.6491 | 1.2175 | 25.739617 | 0.755524 | -0.050722 | 20.459421 | 5.542815 |
-LCM LORA SDXL [28] | 1.6387 | 1.3833 | 27.329910 | 0.8364 | 0.4959 | 21.4824 | 5.657512 |
-Safe SD STRONG [31] | 1.630810 | 1.146611 | 25.576418 | 0.816518 | -0.102223 | 20.621118 | 5.8643 |
-Safe SD MEDIUM [31] | 1.627511 | 1.129815 | 26.279814 | 0.810120 | 0.2042 12 | 20.788012 | 6.1760 |
-Safe SD WEAK [31] | 1.607812 | 1.118817 | 26.118015 | 0.780923 | -0.126424 | 20.387324 | 5.386117 |
-SD v2.1 [40] | 1.552413 | 1.109418 | 26.582312 | 0.8377 | 0.4116 | 21.0502 | 5.307319 |
-SD v2.0 [40] | 1.527714 | 1.130014 | 25.348121 | 0.817017 | 0.0872 18 | 20.752913 | 5.9060 |
-Openjourney v2 [35] | 1.500015 | 0.995620 | 24.698423 | 0.795822 | -0.041521 | 20.408822 | 5.199522 |
-Redshift diffusion [39] | 1.473316 | 1.138212 | 25.157222 | 0.810121 | 0.0218 20 | 20.615519 | 5.7657 |
-Dreamlike Diffusion v1.0 [10] | 1.465217 | 1.2052 | 29.6506 | 0.8543 | 0.6508 | 21.2664 | 5.661411 |
-SD v1.5 [40] | 1.441718 | 1.136213 | 25.497219 | 0.821413 | 0.1686 14 | 20.714316 | 6.0535 |
-IF-I-XL v1.0 [1] | 1.380819 | 0.922122 | 27.4512 | 0.8449 | 0.6087 | 20.747414 | 5.301220 |
-SD v1.4 [40] | 1.359220 | 0.951121 | 25.369720 | 0.819016 | 0.1050 17 | 20.653517 | 5.8571 |
-Vintedois Diffusion v0.1 [51] | 1.356221 | 1.079719 | 26.590111 | 0.8341 | 0.3562 10 | 20.835811 | 5.018123 |
-IF-I-L v1.0 [1] | 1.263522 | 0.881423 | 27.4836 | 0.8384 | 0.4463 | 20.717015 | 5.650213 |
-MultiFusion [29] | 1.237223 | 1.129816 | 23.813324 | 0.815119 | 0.0695 19 | 20.478020 | 4.382424 |
-IF-I-M v1.0 [1] | 1.013524 | 0.792824 | 25.952216 | 0.832911 | 0.1637 15 | 20.403523 | 5.7451 |
-Model | Human | EvalAlign | HPS v2 | CLIP-score | ImageReward | PickScore | IS |
-IF-I-XL v1.0 [1] | 5.4500 | 5.5300 | 32.547710 | 0.8579 | 0.4391 | 21.199810 | 7.1864 |
-IF-I-L v1.0 [1] | 5.2300 | 5.4500 | 32.7140 | 0.8538 | 0.3820 | 21.128412 | 6.657120 |
-SDXL Refiner v1.0 [47] | 5.2100 | 5.4000 | 35.6465 | 0.8528 | 0.4738 | 22.3532 | 7.1101 |
-LCM SDXL [28] | 5.1800 | 5.3300 | 33.8011 | 0.8512 | 0.3833 | 21.9620 | 7.006714 |
-PixArt XL2 1024 MS [6] | 5.1100 | 5.3100 | 37.0493 | 0.8634 | 0.6542 | 22.3926 | 6.929316 |
-IF-I-M v1.0 [1] | 5.0800 | 5.2200 | 31.095114 | 0.8434 | 0.0499 10 | 20.827020 | 7.2586 |
-LCM LORA SDXL [28] | 5.0600 | 5.2700 | 32.7752 | 0.834910 | 0.1618 | 21.7627 | 6.838919 |
-SDXL v1.0 [33] | 5.0300 | 5.3500 | 35.1593 | 0.8540 | 0.4322 | 22.1291 | 7.2762 |
-Wuerstchen [32] | 4.8700 | 5.1700 | 36.4632 | 0.8381 | 0.2513 | 21.7779 | 7.1280 |
-Openjourney [34] | 4.830010 | 4.920015 | 31.149512 | 0.817316 | -0.086714 | 21.116313 | 6.272922 |
-SD v2.1 [40] | 4.800011 | 5.070011 | 31.101713 | 0.827814 | -0.045312 | 21.2093 | 7.0767 |
-MultiFusion [29] | 4.680012 | 4.800018 | 28.795724 | 0.826415 | -0.133715 | 20.962517 | 4.959324 |
-Dreamlike Diffusion v1.0 [10] | 4.660013 | 5.150010 | 34.8196 | 0.8493 | 0.2295 | 21.5550 | 6.858118 |
-SD v2.0 [40] | 4.640014 | 5.010012 | 30.615317 | 0.829813 | -0.142416 | 21.190511 | 7.012413 |
-Vintedois Diffusion v0.1 [51] | 4.620015 | 4.950014 | 31.950311 | 0.831912 | -0.022211 | 21.114114 | 6.198723 |
-Safe SD STRONG [31] | 4.600016 | 4.830017 | 30.661516 | 0.775123 | -0.502822 | 20.749121 | 7.2743 |
-Dreamlike Photoreal v2.0 [11] | 4.560017 | 4.980013 | 33.7712 | 0.834411 | -0.085913 | 21.4832 | 6.936315 |
-Safe SD WEAK [31] | 4.530018 | 4.710020 | 30.564418 | 0.814018 | -0.272818 | 20.989916 | 6.423821 |
-SD v1.4 [40] | 4.520019 | 4.760019 | 29.914920 | 0.804820 | -0.343819 | 20.846219 | 7.015012 |
-SD v1.5 [40] | 4.450020 | 4.900016 | 30.167319 | 0.814217 | -0.221317 | 20.864018 | 7.1642 |
-Safe SD MEDIUM [31] | 4.400021 | 4.560024 | 30.782015 | 0.797421 | -0.359120 | 21.025715 | 7.070910 |
-Redshift diffusion [39] | 4.350022 | 4.670021 | 29.286522 | 0.806619 | -0.417221 | 20.632723 | 7.061011 |
-Safe SD MAX [31] | 4.310023 | 4.590023 | 29.812621 | 0.760124 | -0.609524 | 20.704622 | 7.2273 |
-Openjourney v2 [35] | 4.150024 | 4.650022 | 29.238923 | 0.785122 | -0.605123 | 20.597324 | 6.861317 |
-Table 14 and Table 15 present the comparison results between our proposed method and existing alternatives.
-Owing to the powerful aligned understanding of images and text by MLLMs, our model achieves the best alignment with human performance across all tested benchmarks for 24 text-to-image models, along with a significant improvement in image faithfulness.
-This indicates that our model excels in both comprehending and evaluating the intricate details of generated images, closely mirroring human judgment and setting a new standard for image-text alignment and faithfulness.
-Appendix G Qualitative Analysis
-As shown in Figure 5, we present a comparison of different evaluation metrics on images generated by four models, including human annotated scores, EvalAlign, ImageReward [56], HPSv2 [54], and PickScore [20].
+Appendix F Qualitative Analysis
+As shown in Figure 5, we present a comparison of different evaluation metrics on images generated by four models, including human annotated scores, EvalAlign, ImageReward (Xu et al., 2024), HPSv2 (Wu et al., 2023a), and PickScore (Kirstain et al., 2024).
 The digits in the figure represent the ranking for each evaluation metric, with darker colors indicating higher rankings.
 From the figure, it is evident that our proposed EvalAlign metric closely matches the human rankings across two evaluation dimensions, demonstrating excellent consistency.
-Furthermore, Figure 6 showcases the EvalAlign evaluation metric across different fine-grained results.
-The numbers in the figure represent EvalAlign scores for the corresponding evaluation aspect, with darker colors indicating higher scores and better generation performance.
+Specifically, the numbers in the figure represent EvalAlign scores for the corresponding evaluation aspect, with darker colors indicating higher scores and better generation performance.
 Note that if the text prompt does not specify a particular style, the style consistency score defaults to 0.